### v2

# Invariant subspace perturbations related to defective eigenvalues of $\Delta$-Hermitian and Hamiltonian matrices

Abstract
Structured perturbation results for invariant subspaces of $\Delta$-Hermitian and Hamiltonian matrices are provided. The invariant subspaces under consideration are associated with the eigenvalues perturbed from a single defective eigenvalue. The results show how the original eigenvectors and generalized eigenvectors are involved in composing such perturbed invariant subspaces and eigenvectors.
Keywords perturbation, $\Delta$-Hermitian matrix, Hamiltonian matrix, invariant subspace, defective eigenvalue
AMS subject classification 15A18, 47A55, 65F15
1 Introduction
Eigenvalue problem of Hamiltonian matrices plays a fundamental role in systems and control [4, 7, 16, 19, 24, 29, 30]. For a Hamiltonian matrix, its eigenvalues are symmetric about the imaginary axis and its invariant subspaces normally have special structures as well, e.g., [10, 18]. These structures and symmetry are crucial for solving application problems including optimal and robust controller constructions [2, 3, 11, 16, 19, 30], H2 and H norm computations [5, 6, 25], and distance and robustness problems [1, 8, 9, 14]. In these applications, eigenvalue and invariant subspace perturbation theory plays an important role, in particular those about the imaginary eigenvalues, e.g., [21, 22].
In [22], perturbation results about imaginary eigenvalues were developed for a special type of Hamiltonian matrices with special Hamiltonian perturbations. In this paper we will provide general perturbation results for the eigenvalues of a Hamiltonian matrix that are defective, in the situation that the perturbed matrix is still Hamiltonian. The results are derived based on the perturbation theory for general matrices provided in [28]. Since Hamiltonian matrices can be transformed to $\Delta$-Hermitian matrices (which will be defined later), we will first derive the perturbation results for $\Delta$-Hermitian matrices. We then use the results to establish the perturbation theory for Hamiltonian matrices.
The paper is organized as follows. In Section 2, some results from [28] are provided, which are needed for establishing the structured perturbation theory. Definitions, structured canonical forms, and basic properties of $\Delta$-Hermitian and Hamiltonian matrices are given as well. The perturbation results and derivations for $\Delta$-Hermitian matrices are given in Section 3 and the results for Hamiltonian matrices are in Section 4. Our conclusions are contained in Section 5.
Throughout the paper, $\mathbb{R}$$\mathbb{C}$} represent the real and complex fields, respectively. ${\mathbb{C}}^{m\times n}$ is the space of $m\times n$ complex matrices, and ${\mathbb{C}}^{n}$ is the $n$-dimensional complex column vector space. For a matrix $B$, $B^{\ast}$ and $B^{T}$ are the adjoint (conjugate transpose) and transpose of $B$, respectively. For a square matrix $A$, $\Lambda(A)$ is the spectrum of $A$. We denote by $I_{k}$ and $0_{k}$ the $k\times k$ identity and zero matrices, respectively. For a positive number $d$, the notation $O(t^{d})$ means a term (or matrix) that converges to 0 at the same rate as $t^{d}$ and $o(t^{d})$ means a term (or matrix) that converges faster than $t^{d}$, both when $t\to 0$. We also use notations as $O\left(\left[\begin{array}[]{ccc}t^{d_{11}}&\ldots&t^{d_{1q}}\\ \vdots&\vdots&\vdots\\ t^{d_{p1}}&\ldots&t^{d_{pq}}\end{array}\right]\right)$ to show the orders of different blocks (entries) in a matrix. $t\in{\mathbb{R}}$ is always a positive and sufficiently small parameter. For a given $r$-dimensional subspace $V\subseteq{\mathbb{C}}^{n}$, we call $X\in{\mathbb{C}}^{n\times r}$ a basis matrix of $V$ if $\mathop{\rm range}\nolimits X=V$. Finally, $\imath=\sqrt{-1}$.
2 Preliminaries
Consider the invariant subspaces of the matrix $\lambda I_{p}+N+tB(t)$, where $B(t)=B+O(t)$ for some constant matrix $B$,
$N:=N(m)=\mathop{\rm diag}\nolimits(N_{1},N_{2},\ldots,N_{m}),$ (1)
and
$N_{j}=\left[\begin{array}[]{cccc}0&I_{s_{j}}&&\\ &0&\ddots&\\ &&\ddots&I_{s_{j}}\\ &&&0\end{array}\right]_{j\times j},\quad j=1,\ldots,m$
with integers $s_{1},\ldots,s_{m-1}\geq 0$, $s_{m}>0$, and
$p=s_{1}+2s_{2}+\ldots+ms_{m}.$
By convention, $N_{1}=0_{s_{1}}$. This indicates that $\lambda I_{p}+N$ has $s_{j}$ Jordan blocks of size $j\times j$ for $j=1,\ldots,m$. Because of the term $tB(t)$, the original eigenvalue $\lambda$ will be perturbed with magnitudes of fractional orders of $t$ that depend on the sizes of the Jordan blocks.
We focus on the perturbations of the eigenvalues corresponding to a particular size of Jordan blocks and the corresponding invariant subspaces.
Partition
$B=\left[\begin{array}[]{ccc}B_{11}&\ldots&B_{1m}\\ \vdots&\ddots&\vdots\\ B_{m1}&\ldots&B_{mm}\end{array}\right],\quad B_{ij}=\left[\begin{array}[]{ccc}B_{11}^{(ij)}&\ldots&B_{1j}^{(ij)}\\ \vdots&\ddots&\vdots\\ B_{i1}^{(ij)}&\ldots&B_{ij}^{(ij)}\end{array}\right]$
conformably with the block forms of $N$ and $N_{i},N_{j}$, respectively, for $1\leq i,j\leq m$. Let $W_{m}=B_{m1}^{(mm)}$ and for any $k\in\{1,\ldots,m-1\}$, define
$W_{k}=\left[\begin{array}[]{c|ccc}B_{k1}^{(kk)}&B_{k1}^{(k,k+1)}&\ldots&B_{k1}^{(km)}\\ \hline\cr B_{k+1,1}^{(k+1,k)}&B_{k+1,1}^{(k+1,k+1)}&\ldots&B_{k+1,1}^{(k+1,m)}\\ \vdots&\vdots&\ddots&\vdots\\ B_{m1}^{(mk)}&B_{m1}^{(m,k+1)}&\ldots&B_{m1}^{(mm)}\end{array}\right]=\left[\begin{array}[]{c|c}B_{k1}^{(kk)}&W_{k,k+1}\\ \hline\cr W_{k+1,k}&W_{k+1}\end{array}\right].$ (2)
The following perturbation results are given in [17, 23, 28].
Lemma 1
Consider the matrix $\lambda I_{p}+N+tB(t)$ with $B(t)=B+O(t)$ and $N:=N(m)$. Let $W_{k}$ be defined in (2). Suppose $\rho\in\{1,\ldots,m\}$ and $W_{\rho+1}$ is invertible when $\rho<m$. Let $\gamma_{1}^{(\rho)},\ldots,\gamma_{s_{\rho}}^{(\rho)}$ be the $s_{\rho}$ eigenvalues of the matrix
$S_{\rho}=\left\{\begin{array}[]{ll}B_{\rho 1}^{(\rho\rho)}+W_{\rho,\rho+1}G_{\rho}&\rho<m\\ B_{m1}^{(mm)}&\rho=m,\end{array}\right.,\qquad G_{\rho}=-W_{\rho+1}^{-1}W_{\rho+1,\rho}=:\left[\begin{array}[]{c}G_{\rho+1,\rho}\\ \vdots\\ G_{m\rho}\end{array}\right].$ (3)
Assume $t$ is sufficiently small.
  1. (a)
[…]
    $\lambda_{ij}^{(\rho)}(t)=\lambda+t^{1/\rho}\mu_{ij}^{(\rho)}+o(t^{1/\rho}),\quad i=1,\ldots,s_{\rho},\quad j=1,\ldots,\rho.$
  2. (b)
    Suppose all $W_{1},\ldots,W_{m}$ are invertible. Let
    $\Theta_{\rho}(S_{\rho})=\left[\begin{array}[]{cccc}0&I_{s_{\rho}}&&\\ &0&\ddots&\\ &&\ddots&I_{s_{\rho}}\\ S_{\rho}&&&0\end{array}\right]_{\rho\times\rho},$ (4)
    where $S_{\rho}$ is defined in (3). Then there is a matrix
    $X_{\rho}(t)=\left[\begin{array}[]{ccccc}X_{1\rho}^{T}(t)&\ldots&X_{\rho\rho}^{T}(t)&\ldots&X_{m\rho}^{T}(t)\end{array}\right]^{T}+O(t),$ (5)
    $X_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cc}I_{s_{\rho}}&0\\ 0&0\end{array}\right]+\mathop{\rm diag}\nolimits(0_{s_{\rho}},t^{1/\rho}I_{s_{\rho}},\ldots,t^{1-1/\rho}I_{s_{\rho}})&k=\rho\\ \left[\begin{array}[]{cc}G_{k\rho}&0\\ 0&0\end{array}\right]+O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho.\end{array}\right$.
    where $G_{\rho+1,\rho},\ldots,G_{m\rho}$ are given in (3), such that
    $(\lambda I_{p}+N+tB(t))X_{\rho}(t)=X_{\rho}(t)(\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t))$
    for some $\widetilde{\Theta}_{\rho}(t)=\Theta_{\rho}(S_{\rho})+O(t^{1/\rho})$.
Proof. The results can be derived by applying those in [17, 23, 28] to the special case when the original matrix has a single eigenvalue $\lambda$ and is already in the Jordan canonical form.
We point out that $S_{\rho}$ is the Schur complement of $W_{\rho}$ corresponding to $W_{\rho+1}$. Note $\Lambda(\Theta_{\rho}(S_{\rho}))=\{\mu^{(\rho)}_{ij}\}$. Note also if $W_{1},\ldots,W_{m}$ are invertible, $S_{\rho}$ is invertible for any $\rho$. Then $\Theta_{\rho}(S_{\rho})$ is invertible as well, and all $\mu_{ij}^{(\rho)}$, $1\leq i\leq s_{\rho}$, $1\leq j\leq\rho$, are nonzero.
Remark 1
The assumption that $W_{1},\ldots,W_{m}$ are invertible in Lemma 1(b) is generic, and is necessary for deriving the invariant subspace perturbation results, e.g., [28].
The following result shows the perturbation behaviors of the invariant subspaces corresponding to a subset of $\Lambda(\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t))$. Suppose $F\in{\mathbb{C}}^{n\times\ell}$ is a full column rank matrix and
$\Theta_{\rho}(S_{\rho})F=F\Omega,$
where $S_{\rho}$ and $\Theta_{\rho}(S_{\rho})$ are given in (3) and (4), and $\Omega\in{\mathbb{C}}^{\ell\times\ell}$. Clearly $\Lambda(\Omega)\subseteq\Lambda(\Theta_{\rho}(S_{\rho}))$. Based on the block structure of $\Theta_{\rho}(S_{\rho})$, we have the expression
$F=\left[\begin{array}[]{cccc}Q^{T}&(Q\Omega)^{T}&\ldots&(Q\Omega^{\rho-1})^{T}\end{array}\right]^{T}$ (6)
for some matrix $Q\in{\mathbb{C}}^{s_{\rho}\times\ell}$ that satisfies
$S_{\rho}Q=Q\Omega^{\rho}.$
Lemma 2
Assume the conditions of Lemma 1 hold and $\Lambda(\Omega)\cap(\Lambda(\Theta_{\rho}(S_{\rho}))\backslash\Lambda(\Omega))=\emptyset.$ Then
$(\lambda I_{p}+N+tB(t))X_{\Omega}(t)=X_{\Omega}(t)(\lambda I_{\ell}+t^{1/\rho}\widetilde{\Omega}(t)),$
where $\widetilde{\Omega}(t)=\Omega+O(t^{1/\rho})$,
$X_{\Omega}(t)=\left[\begin{array}[]{ccccc}X_{\Omega,1}^{T}(t)&\ldots&X_{\Omega,\rho}^{T}(t)&\ldots&X_{\Omega,m}^{T}(t)\end{array}\right]^{T}+O(t),$
$X_{\Omega,k}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}Q^{T}&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}&\\ \qquad\qquad\qquad\qquad\qquad\quad+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ \left[\begin{array}[]{cccc}(G_{k\rho}Q)^{T}&0&\ldots&0\end{array}\right]^{T}&\\ \qquad+O(\left[\begin{array}[]{cccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t^{1-1/\rho}&t&\ldots&t\end{array}\right]^{T})&k>\rho,\end{array}\right$.
and $Q$ is given in (6).
Proof. See [28].
Let $\Delta$ be Hermitian and invertible. A matrix $C$ is $\Delta$-Hermitian if
$\Delta C=(\Delta C)^{\ast}=C^{\ast}\Delta.$
The structured Jordan canonical form of a $\Delta$-Hermitian is well known.
Proposition 3
Suppose $C$ is $\Delta$-Hermitian. There exists an invertible matrix $U$ such that
$U^{-1}CU$ $=$ $\mathop{\rm diag}\nolimits(C_{1},\ldots,C_{r},C_{r+1},\ldots,C_{q})=:\widehat{C}$,
$U^{\ast}\Delta U$ $=$ $\mathop{\rm diag}\nolimits(\Delta_{1},\ldots,\Delta_{r},\Delta_{r+1},\ldots,\Delta_{q})=:\widehat{\Delta}=\widehat{\Delta}^{\ast}=\widehat{\Delta}^{-1}.$
  1. (i)
    For each $j=1,\ldots,r$, $C_{j}$ contains a pair of complex conjugate eigenvalues $\lambda_{j},\bar{\lambda}_{j}$,
    $C_{j}=\left[\begin{array}[]{cc}\lambda_{j}I_{p_{j}}+N(m_{j})&0\\ 0&(\lambda_{j}I_{p_{j}}+N(m_{j}))^{\ast}\end{array}\right],\quad\Delta_{j}=\left[\begin{array}[]{cc}0&I_{p_{j}}\\ I_{p_{j}}&0\end{array}\right]=\Delta_{j}^{\ast}=\Delta_{j}^{-1},$
    where $N(m_{j})$ is defined in (1) with $s^{(j)}_{k}$ size $k\times k$ Jordan blocks for $k=1,\ldots,m_{j}$, and $p_{j}=\sum_{k=1}^{m_{j}}ks^{(j)}_{k}$.
    Moreover, $\lambda_{1},\ldots,\lambda_{r}$ are distinct.
  2. (ii)
    For each $j=r+1,\ldots,q$, $C_{j}$ contains a real eigenvalue $\alpha_{j}$,
    $C_{j}=\alpha_{j}I_{p_{j}}+N(m_{j}),\quad\Delta_{j}=\mathop{\rm diag}\nolimits(\Gamma_{1}^{(j)},\Gamma_{2}^{(j)},\ldots,\Gamma_{m_{j}}^{(j)})=\Delta^{\ast}_{j}=\Delta_{j}^{-1},$
    and
[…]
    $\Sigma_{k}^{(j)}=\left[\begin{array}[]{cc}I_{t^{(j)}_{k}}&0\\ 0&-I_{s^{(j)}_{k}-t^{(j)}_{k}}\end{array}\right],\qquad k=1,\ldots,m_{j},$
    where $N(m_{j})$ and $p_{j}$ are defined in the same way as in (i) and $t_{k}^{(j)}\geq 0$ for $k=1,\ldots,m_{j}$. The diagonal entries of $\Sigma_{k}^{(j)}$ (which are $\pm 1$) are the sign signatures of $C$ corresponding to the $k\times k$ Jordan blocks of $\alpha_{j}$.
    Moreover, $\alpha_{r+1},\ldots,\alpha_{q}$ are distinct.
Proof. This is a slightly modified form of those in [12, 20, 27].
Define $J_{n}=\left[\begin{array}[]{cc}0&I_{n}\\ -I_{n}&0\end{array}\right]$. A matrix $H\in{\mathbb{C}}^{2n\times 2n}$ is Hamiltonian if $J_{n}H=(J_{n}H)^{\ast}$.
Proposition 4
Suppose $H$ is Hamiltonian. There exists an invertible matrix $U$ such that
$U^{-1}HU=\mathop{\rm diag}\nolimits(H_{1},\ldots,H_{r},H_{r+1},\ldots,H_{q}),\quad U^{\ast}J_{n}U=\mathop{\rm diag}\nolimits(J_{p_{1}},\ldots,J_{p_{r}},\imath\Delta_{r+1},\ldots,\imath\Delta_{q}).$
  1. (i)
    For each $j=1,\ldots,r$, $H_{j}$ contains a pair of nonimaginary eigenvalues $\lambda_{j},-\bar{\lambda}_{j}$,
    $H_{j}=\left[\begin{array}[]{cc}\lambda_{j}I_{p_{j}}+N(m_{j})&0\\ 0&-(\lambda_{j}I_{p_{j}}+N(m_{j}))^{\ast}\end{array}\right],$
    where $N(m_{j})$ and $p_{j}$ are as defined in Proposition 3(i).
    Moreover, $\lambda_{1},\ldots,\lambda_{r}$ are distinct.
  2. (ii)
    For each $j=r+1,\ldots,q$, $H_{j}$ contains an imaginary eigenvalue $\imath\alpha_{j}$,
    $H_{j}=\imath(\alpha_{j}I_{p_{j}}+N(m_{j})),\quad\Delta_{j}=\mathop{\rm diag}\nolimits(\Gamma_{1}^{(j)},\Gamma_{2}^{(j)},\ldots,\Gamma^{(j)}_{m_{j}})$,
    where $N(m_{j})$ and $\Delta_{j}$ are the same as that in Proposition 3(ii). The diagonal entries of $\Sigma_{k}^{(j)}$ in $\Gamma_{k}^{(j)}$ are the structure inertia indices of $H$ corresponding to the $k\times k$ Jordan blocks of $\imath\alpha_{j}$.
    Moreover, $\imath\alpha_{r+1},\ldots,\imath\alpha_{q}$ are distinct.
Proof. This is a slightly modified form of those in [10, 15, 18].
3 Invariant subspace perturbations of $\Delta$-Hermitian matrices
We investigate how the invariant subspaces change when a $\Delta$-Hermitian matrix $C$ is perturbed to $C+tD$, where $D$ is another given $\Delta$-Hermitian matrix and $t$ is sufficiently small. For an invariant subspace corresponding to certain eigenvalues of $C$ that are different from the rest, the perturbation results are well known, e.g., [26]. Our focus is about the invariant subspaces of $C+tD$ corresponding to the eigenvalues that are perturbed from a single defective eigenvalue of $C$ with the same fractional order. We consider
  1. (a)
    a single defective nonreal eigenvalue $\lambda$ (together with $\bar{\lambda}$ due to the pairing)
  2. (b)
    a single defective real eigenvalue $\alpha$.
From Proposition 3, there is an invertible matrix $U=\left[\begin{array}[]{cc}U_{T}&U_{C}\end{array}\right]$ such that
$CU=U\left[\begin{array}[]{cc}C_{T}&0\\ 0&C_{C}\end{array}\right],\quad U^{\ast}\Delta U=\left[\begin{array}[]{cc}\Delta_{T}&0\\ 0&\Delta_{C}\end{array}\right].$ (7)
In case (a),
$C_{T}=\left[\begin{array}[]{cc}\lambda I_{p}+N&0\\ 0&(\lambda I_{p}+N)^{\ast}\end{array}\right],\quad\Delta_{T}=\left[\begin{array}[]{cc}0&I_{p}\\ I_{p}&0\end{array}\right],$ (8)
where $\lambda$ is nonreal and $\lambda,\bar{\lambda}\notin\Lambda(C_{C})$, $N:=N(m)$ is defined in (1) and $p=\sum_{k=1}^{m}ks_{k}$. In case (b),
$C_{T}=\alpha I_{p}+N,\quad\Delta_{T}=\mathop{\rm diag}\nolimits(\Gamma_{1},\ldots,\Gamma_{m}),$ (9)
[…]
$\tilde{s}=\left\{\begin{array}[]{ll}\sum_{j=0}^{q-1}(t_{2j+1}+js_{2j+1})+\sum_{j=1}^{q}js_{2j}&m=2q\\ \sum_{j=0}^{q}(t_{2j+1}+js_{2j+1})+\sum_{j=1}^{q}js_{2j}&m=2q+1\end{array}\right.\qquad\tilde{t}=p-\tilde{s}$. (10)
From the second relation in (7),
$U^{-1}=\left[\begin{array}[]{cc}\Delta_{T}&0\\ 0&\Delta_{C}\end{array}\right]^{-1}U^{\ast}\Delta.$
Define
$\left[\begin{array}[]{cc}D_{11}&D_{12}\\ D_{21}&D_{22}\end{array}\right]:=U^{-1}DU=\left[\begin{array}[]{cc}\Delta_{T}&0\\ 0&\Delta_{C}\end{array}\right]^{-1}U^{\ast}\Delta DU.$ (11)
One has
$U^{-1}(C+tD)U=\left[\begin{array}[]{cc}C_{T}+tD_{11}&tD_{12}\\ tD_{21}&C_{C}+tD_{22}\end{array}\right],$
which is $\mathop{\rm diag}\nolimits(\Delta_{T},\Delta_{C})$-Hermitian.
Consider the Riccati equation
$(C_{C}+tD_{22})Y(t)-Y(t)(C_{T}+tD_{11})-tY(t)D_{12}Y(t)+tD_{21}=0.$ (12)
Since $\Lambda(C_{C})\cap\Lambda(C_{T})=\emptyset$, when $t$ is sufficiently small it has a solution $Y(t)=O(t).$ Note in both cases, $\Delta_{T}^{-1}=\Delta_{T}$. By taking the adjoint on both sides of (12) and using the fact that $U^{-1}(C+tD)U$ is $\mathop{\rm diag}\nolimits(\Delta_{T},\Delta_{C})$-Hermitian, one has
$(C_{T}+tD_{11})\Delta_{T}Y^{\ast}(t)\Delta_{C}-\Delta_{T}Y^{\ast}(t)\Delta_{C}(C_{C}+tD_{22})-tD_{12}+t\Delta_{T}Y^{\ast}(t)\Delta_{C}D_{21}\Delta_{T}Y^{\ast}(t)\Delta_{C}=0.$
Let
$\widehat{Z}(t)=\left[\begin{array}[]{cc}I&-\Delta_{T}Y^{\ast}(t)\Delta_{C}\\ Y(t)&I\end{array}\right].$
One has
$(C+tD)U\widehat{Z}(t)=U\widehat{Z}(t)\left[\begin{array}[]{cc}\widehat{C}_{T}(t)&0\\ 0&\widehat{C}_{C}(t)\end{array}\right],\quad(U\widehat{Z}(t))^{\ast}\Delta(U\widehat{Z}(t))=\left[\begin{array}[]{cc}\widehat{\Delta}_{T}(t)&0\\ 0&\widehat{\Delta}_{C}(t)\end{array}\right],$
where
$\widehat{C}_{T}(t)=C_{T}+tD_{11}+tD_{12}Y(t),\qquad\widehat{C}_{C}(t)=C_{C}+tD_{22}-tD_{21}\Delta_{T}Y^{\ast}(t)\Delta_{C},$
and by using $\Delta_{T}^{\ast}=\Delta_{T}=\Delta_{T}^{-1}$, $\Delta_{C}^{\ast}=\Delta_{C}$,
$\widehat{\Delta}_{T}(t)=\Delta_{T}+Y^{\ast}(t)\Delta_{C}Y(t),\quad\widehat{\Delta}_{C}(t)=\Delta_{C}+\Delta_{C}Y(t)\Delta_{T}Y^{\ast}(t)\Delta_{C}.$
It is straightforward to verify
$\widehat{Z}^{-1}(t)=\left[\begin{array}[]{cc}\widehat{\Delta}_{T}(t)&0\\ 0&\widehat{\Delta}_{C}(t)\end{array}\right]^{-1}\left[\begin{array}[]{cc}\Delta_{T}&0\\ 0&\Delta_{C}\end{array}\right]\left[\begin{array}[]{cc}I&\Delta_{T}Y^{\ast}(t)\Delta_{C}\\ -Y(t)&I\end{array}\right].$ (13)
We now show that the matrix $\widehat{\Delta}_{T}(t)$ is congruent to $\Delta_{T}$ with a matrix close to $I$. For $\Delta_{T}$ there is a unitary matrix $Q$ such that
[…]
where $\tilde{s}$ and $\tilde{t}$ are both $p$ in case (a) and are given in (10) in case (b). Partition
$Q^{\ast}Y^{\ast}(t)\Delta_{C}Y(t)Q=\left[\begin{array}[]{cc}Y_{11}(t)&Y_{12}(t)\\ Y_{12}^{\ast}(t)&Y_{22}(t)\end{array}\right]=O(t^{2}),$
where $Y_{11}(t)=Y_{11}^{\ast}(t)\in{\mathbb{C}}^{\tilde{s}\times\tilde{s}}$, $Y_{22}(t)=Y_{22}^{\ast}(t)\in{\mathbb{C}}^{\tilde{t}\times\tilde{t}}$. When $t$ is sufficiently small, the matrix
$Z_{12}(t)=-(I_{\tilde{s}}+Y_{11}(t))^{-1}Y_{12}(t)=O(t^{2})$
is well defined, and there are matrices $Z_{11}(t)=I_{\tilde{s}}+O(t^{2})$ and $Z_{22}(t)=I_{\tilde{t}}+O(t^{2})$ such that
$Z_{11}^{\ast}(t)Z_{11}(t)$ $=$ $I_{\tilde{s}}+Y_{11}(t)$
$Z_{22}^{\ast}(t)Z_{22}(t)$ $=$ $I_{\tilde{t}}-Y_{22}(t)+Y_{12}^{\ast}(t)(I_{\tilde{s}}+Y_{11}(t))^{-1}Y_{12}(t).$
Define
$Z_{1}(t)=Q\left[\begin{array}[]{cc}I_{\tilde{s}}&Z_{12}(t)\\ 0&I_{\tilde{t}}\end{array}\right]\left[\begin{array}[]{cc}Z_{11}^{-1}(t)&0\\ 0&Z_{22}^{-1}(t)\end{array}\right]Q^{\ast}=I+O(t^{2}).$
One has
$Z_{1}^{\ast}(t)\widehat{\Delta}_{T}(t)Z_{1}(t)=\Delta_{T}.$ (14)
Similarly, there is a $Z_{2}(t)=I+O(t^{2})$ such that
$Z_{2}^{\ast}(t)\widehat{\Delta}_{C}(t)Z_{2}(t)=\Delta_{C}.$ (15)
Define
$Z(t)=\widehat{Z}(t)\left[\begin{array}[]{cc}Z_{1}(t)&0\\ 0&Z_{2}(t)\end{array}\right].$
Then from (13), (14), and (15), one has
$Z^{-1}(t)=\left[\begin{array}[]{cc}\Delta_{T}Z^{\ast}_{1}(t)\Delta_{T}&0\\ 0&\Delta_{C}^{-1}Z_{2}^{\ast}(t)\Delta_{C}\end{array}\right]\left[\begin{array}[]{cc}I&\Delta_{T}Y^{\ast}(t)\Delta_{C}\\ -Y(t)&I\end{array}\right],$
and
$(C+tD)UZ(t)=UZ(t)\left[\begin{array}[]{cc}\widetilde{C}_{T}(t)&0\\ 0&\widetilde{C}_{C}(t)\end{array}\right]$ (18)
$\widetilde{C}_{T}(t)=Z_{1}^{-1}(t)\widehat{C}_{T}(t)Z_{1}(t)=Z_{1}^{-1}(t)(C_{T}+tD_{11}+tD_{12}Y(t))Z_{1}(t)$
$\widetilde{C}_{C}(t)=Z_{2}^{-1}(t)\widehat{C}_{C}(t)Z_{2}(t)=Z_{2}^{-1}(t)(C_{C}+tD_{22}-tD_{21}\Delta_{T}Y^{\ast}(t)\Delta_{C})Z_{2}(t)$
$(UZ(t))^{\ast}\Delta(UZ(t))=Z^{\ast}(t)\left[\begin{array}[]{cc}\Delta_{T}&0\\ 0&\Delta_{C}\end{array}\right]Z(t)=\left[\begin{array}[]{cc}\Delta_{T}&0\\ 0&\Delta_{C}\end{array}\right].$ (24)
Obviously, $\mathop{\rm diag}\nolimits(\widetilde{C}_{T}(t),\widetilde{C}_{C}(t))$ is $\mathop{\rm diag}\nolimits(\Delta_{T},\Delta_{C})$-Hermitian. Note that the columns of
$U\left[\begin{array}[]{c}I\\ Y(t)\end{array}\right]Z_{1}(t),\qquad U\left[\begin{array}[]{c}-\Delta_{T}Y^{\ast}(t)\Delta_{C}\\ I\end{array}\right]Z_{2}(t)$
span the invariant subspaces of $C+tD$ corresponding to the eigenvalues of $\widetilde{C}_{T}(t)$ and $\widetilde{C}_{C}(t)$, respectively.
In the following, we consider the perturbations of the invariant subspaces corresponding to the eigenvalues perturbed from a single defective eigenvalue of $C$ of the same fractional order.
3.1 Nonreal eigenvalue case
In this case $C_{T}$ and $\Delta_{T}$ are given in (8). We need to transform $\widetilde{C}_{T}(t)$ further to a block diagonal form as $C_{T}$. Because $Z_{1}(t)=I+O(t^{2})$ and $Y(t)=O(t)$, we have the expression
$\widetilde{C}_{T}(t)=C_{T}+t\widetilde{D}_{11}(t),\qquad\widetilde{D}_{11}(t)=D_{11}+O(t).$
Recall $\widetilde{C}_{T}(t)$ and $C_{T}$ are $\Delta_{T}$-Hermitian. So is $\widetilde{D}_{11}(t)$. In this case $\Delta_{T}=\left[\begin{array}[]{cc}0&I_{p}\\ I_{p}&0\end{array}\right]$. Partition
$D_{11}=\left[\begin{array}[]{cc}M_{11}&M_{12}\\ M_{21}&M_{11}^{\ast}\end{array}\right],\qquad\widetilde{D}_{11}(t)=\left[\begin{array}[]{cc}\widetilde{M}_{11}(t)&\widetilde{M}_{12}(t)\\ \widetilde{M}_{21}(t)&\widetilde{M}_{11}^{\ast}(t)\end{array}\right],$
conformably, where $M_{12}=M_{12}^{\ast}$, $M_{21}=M_{21}^{\ast}$, $\widetilde{M}_{12}(t)=\widetilde{M}_{12}^{\ast}(t)$, and $\widetilde{M}_{21}(t)=\widetilde{M}_{21}^{\ast}(t)$. As before, we perform a structure preserving transformation on $\widetilde{C}_{T}(t)$ to eliminate the (1,2) and (2,1) blocks. Since $\lambda\neq\bar{\lambda}$, when $t$ is sufficiently small, the Riccati equation
$(\lambda I_{p}+N+t\widetilde{M}_{11}(t))^{\ast}\widetilde{Z}_{21}(t)-\widetilde{Z}_{21}(t)(\lambda I_{p}+N+t\widetilde{M}_{11}(t))-t\widetilde{Z}_{21}(t)\widetilde{M}_{12}(t)\widetilde{Z}_{21}(t)+t\widetilde{M}_{21}(t)=0$
has a unique skew-Hermitian solution $\widetilde{Z}_{21}(t)=-\widetilde{Z}_{21}^{\ast}(t)=O(t)$. Let $\widetilde{Z}_{12}(t)=-\widetilde{Z}_{12}^{\ast}(t)=O(t)$ be the unique skew-Hermitian solution of the Sylvester equation
$(\lambda I_{p}+N+t(\widetilde{M}_{11}(t)+\widetilde{M}_{12}(t)\widetilde{Z}_{21}(t)))\widetilde{Z}_{12}(t)-\widetilde{Z}_{12}(t)(\lambda I_{p}+N+t(\widetilde{M}_{11}(t)+\widetilde{M}_{12}(t)\widetilde{Z}_{21}(t)))^{\ast}$
$+t\widetilde{M}_{12}(t)=0.$
Then
$\widetilde{Z}_{1}(t)=\left[\begin{array}[]{cc}I_{p}&0\\ \widetilde{Z}_{21}(t)&I_{p}\end{array}\right]\left[\begin{array}[]{cc}I_{p}&\widetilde{Z}_{12}(t)\\ 0&I_{p}\end{array}\right]=\left[\begin{array}[]{cc}I_{p}&\widetilde{Z}_{12}(t)\\ \widetilde{Z}_{21}(t)&I_{p}+\widetilde{Z}_{21}(t)\widetilde{Z}_{12}(t)\end{array}\right]=I_{2p}+O(t)$
satisfies
$\widetilde{Z}_{1}^{\ast}(t)\Delta_{T}\widetilde{Z}_{1}(t)=\Delta_{T}.$ (25)
and
$\widetilde{C}_{T}(t)\widetilde{Z}_{1}(t)=\widetilde{Z}_{1}(t)\left[\begin{array}[]{cc}\widetilde{C}_{1}(t)&0\\ 0&\widetilde{C}_{1}^{\ast}(t)\end{array}\right],\qquad\widetilde{C}_{1}(t)=\lambda I_{p}+N+t(\widetilde{M}_{11}(t)+\widetilde{M}_{12}(t)\widetilde{Z}_{21}(t)).$
Hence,
$(C+tD)UZ(t)\left[\begin{array}[]{cc}\widetilde{Z}_{1}(t)&0\\ 0&I\end{array}\right]=UZ(t)\left[\begin{array}[]{cc}\widetilde{Z}_{1}(t)&0\\ 0&I\end{array}\right]\left[\begin{array}[]{cc|c}\widetilde{C}_{1}(t)&0&0\\ 0&\widetilde{C}_{1}^{\ast}(t)&0\\ \hline\cr 0&0&\widetilde{C}_{C}(t)\end{array}\right]$ (33)
$\left(UZ(t)\left[\begin{array}[]{cc}\widetilde{Z}_{1}(t)&0\\ 0&I\end{array}\right]\right)^{\ast}\Delta UZ(t)\left[\begin{array}[]{cc}\widetilde{Z}_{1}(t)&0\\ 0&I\end{array}\right]=\left[\begin{array}[]{c|c}\Delta_{T}&0\\ \hline\cr 0&\Delta_{C}\end{array}\right].$ (41)
We now consider the invariant subspace perturbations of $\widetilde{C}_{1}(t)$ and $\widetilde{C}_{1}^{\ast}(t)$ corresponding to the eigenvalues perturbed from $\lambda$ and $\bar{\lambda}$, respectively, with the same fractional order. As discussed in [28], the high order terms in $\widetilde{C}_{1}(t)$ will not affect the perturbation results. So we only consider the matrix
$C_{1}(t)=\lambda I_{p}+N+tM_{11},$
where $M_{11}$ is the (1,1) block of $D_{11}$ defined in (11).
Partition $U_{T}$ in $U$ as $U_{T}=\left[\begin{array}[]{cc}V&V_{c}\end{array}\right]$. Following (7),
$CV=V(\lambda I_{p}+N),\quad CV_{c}=V_{c}(\lambda I_{p}+N)^{\ast},\quad\left[\begin{array}[]{c}V^{\ast}\\ V_{c}^{\ast}\end{array}\right]\Delta\left[\begin{array}[]{cc}V&V_{c}\end{array}\right]=\Delta_{T}=\left[\begin{array}[]{cc}0&I_{p}\\ I_{p}&0\end{array}\right].$
Using the fact that $C$ is $\Delta$-Hermitian,
$(\Delta V)^{\ast}C=(\lambda I_{p}+N)^{\ast}(\Delta V)^{\ast},\quad(\Delta V_{c})^{\ast}C=(\lambda I_{p}+N)(\Delta V_{c})^{\ast}.$
Partition
$V=\left[\begin{array}[]{cccc}V_{1}&V_{2}&\ldots&V_{m}\end{array}\right],\quad V_{c}=\left[\begin{array}[]{cccc}V_{1}^{(c)}&V_{2}^{(c)}&\ldots&V_{m}^{(c)}\end{array}\right],$
with the column numbers of $V_{j},V_{j}^{(c)}$ being identical to that of $N_{j}$, for $j=1,\ldots,m$. For each $V_{j}$ and $V_{j}^{(c)}$, partition further
$V_{j}=\left[\begin{array}[]{cccc}V_{j1}&V_{j2}&\ldots&V_{jj}\end{array}\right],\qquad V_{j}^{(c)}=\left[\begin{array}[]{cccc}V_{j1}^{(c)}&V_{j2}^{(c)}&\ldots&V_{jj}^{(c)}\end{array}\right],$ (42)
with the column numbers of $V_{ji},V_{ji}^{(c)}$ being $s_{j}$ for $i=1,\ldots,j$ and $j=1,\ldots,m$. Define
$E_{j}=\left[\begin{array}[]{cccc}V_{j1}&V_{j+1,1}&\ldots&V_{m1}\end{array}\right],\quad E_{j}^{(c)}=\left[\begin{array}[]{cccc}V_{jj}^{(c)}&V_{j+1,j+1}^{(c)}&\ldots&V_{mm}^{(c)}\end{array}\right],$ (43)
for $j=1,\ldots,m$. Note $E_{1}$ is a basis matrix of the right eigenvector space of $C$ corresponding to the eigenvalue $\lambda$ and $E_{1}^{(c)}$ is a basis matrix of the right eigenvector space of $C$ corresponding to $\bar{\lambda}$. $\Delta E_{1}^{(c)}$ and $\Delta E_{1}$ are basis matrices of the left eigenvector spaces of $C$ corresponding to the eigenvalue $\lambda$ and $\bar{\lambda}$, respectively. From (11),
$D_{11}=\Delta_{T}U_{T}^{\ast}\Delta DU_{T}.$
Then
$M_{11}=V_{c}^{\ast}\Delta DV.$ (44)
Let the matrix $B$ in Lemma 1 be $M_{11}$. Then following (2), $W_{m}=(V_{mm}^{(c)})^{\ast}\Delta DV_{m1}$ and
$W_{k}=(E_{k}^{(c)})^{\ast}(\Delta D)E_{k}=\left[\begin{array}[]{cc}(V_{kk}^{(c)})^{\ast}\Delta DV_{k1}&(V_{kk}^{(c)})^{\ast}\Delta DE_{k+1}\\ (E_{k+1}^{(c)})^{\ast}\Delta DV_{k1}&W_{k+1}\end{array}\right]$ (45)
for $k=1,\ldots,m-1$. Similarly, for $\widetilde{C}_{1}^{\ast}(t)$ we consider the matrix
$C_{1}^{\ast}(t)=(\lambda I_{p}+N+tM_{11})^{\ast}=\bar{\lambda}I_{p}+N^{\ast}+tM_{11}^{\ast}.$
Let
[…]
Then
$P^{\ast}C_{1}^{\ast}(t)P=\bar{\lambda}I_{p}+N+tP^{\ast}M_{11}^{\ast}P.$ (47)
Since $P^{\ast}M_{11}^{\ast}P=P^{\ast}V^{\ast}(\Delta D)V_{c}P$, the matrix corresponding to $W_{k}$ in (2) is
$\widetilde{W}_{k}=E_{k}^{\ast}(\Delta D)E_{k}^{(c)}=W_{k}^{\ast},$ (48)
for $k=1,\ldots,m$. We have the following results.
Theorem 5
Let $\Delta$ be Hermitian and invertible. Suppose both $C$ and $D$ are $\Delta$-Hermitian matrices and $C+tD$ has the structured decomposition (34) when $t$ is sufficiently small. Let $W_{1},\ldots,W_{m}$ be defined in (45) corresponding to the eigenvalue $\lambda$ of $C$. For a fixed $\rho\in\{1,\ldots,m\}$, assume $W_{\rho+1}$ is invertible if $\rho<m$. Let
$G_{\rho}=-W_{\rho+1}^{-1}(E_{\rho+1}^{(c)})^{\ast}\Delta DV_{\rho 1},\qquad G^{(c)}_{\rho}=-W_{\rho+1}^{-\ast}E_{\rho+1}^{\ast}\Delta DV_{\rho\rho}^{(c)},$
and let $\gamma_{1}^{(\rho)},\ldots,\gamma_{s_{\rho}}^{(\rho)}$ be the $s_{\rho}$ eigenvalues of the matrix
$S_{\rho}=\left\{\begin{array}[]{ll}(V_{\rho\rho}^{(c)})^{\ast}\Delta D(V_{\rho 1}+E_{\rho+1}G_{\rho})&\rho<m\\ (V_{mm}^{(c)})^{\ast}\Delta DV_{m1}&\rho=m,\end{array}\right$.
where $E_{\rho+1}$, $E_{\rho+1}^{(c)}$, $V_{\rho 1}$, $V_{\rho\rho}^{(c)}$ are defined in (43) and (42), respectively.
  1. (a)
[…]
    $\lambda_{ij}^{(\rho)}(t)=\lambda+t^{1/\rho}\mu_{ij}^{(\rho)}+o(t^{1/\rho}),\quad i=1,\ldots,s_{\rho},\quad j=1,\ldots,\rho.$
  2. (b)
    Suppose all $W_{1},\ldots,W_{m}$ are invertible. Let $\Theta_{\rho}(S_{\rho})$ be defined as in (4) with $S_{\rho}$ given above. There are matrices
    $\widetilde{V}_{\rho}(t)$ $=$ $E_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}&0\\ G_{\rho}&0\end{array}\right]+\sum_{k=1}^{m}V_{k}\widetilde{X}_{k\rho}(t)+O(t),$
    $\widetilde{V}_{\rho}^{(c)}(t)$ $=$ $E_{\rho}^{(c)}\left[\begin{array}[]{cc}0&I_{s_{\rho}}\\ 0&G_{\rho}^{(c)}\end{array}\right]+\sum_{k=1}^{m}V_{k}^{(c)}\widetilde{X}_{k\rho}^{(c)}(t)+O(t),$
    with
    $\widetilde{X}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \mathop{\rm diag}\nolimits(0_{s_{\rho}},t^{1/\rho}I_{s_{\rho}},\ldots,t^{1-1/\rho}I_{s_{\rho}})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
    and
    $\widetilde{X}^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \mathop{\rm diag}\nolimits(t^{1-1/\rho}I_{s_{\rho}},\ldots,t^{1/\rho}I_{s_{\rho}},0_{s_{\rho}})+O(\left[\begin{array}[]{cccc}t&t^{1-1/\rho}&\ldots&t^{1/\rho}\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho,\end{array}\right$.
    such that the matrix $\widetilde{U}_{\rho}(t)=\left[\begin{array}[]{cc}\widetilde{V}_{\rho}(t)&\widetilde{V}_{\rho}^{(c)}(t)\end{array}\right]$ satisfies
    $(C+tD)\widetilde{U}_{\rho}(t)=\widetilde{U}_{\rho}(t)\left[\begin{array}[]{cc}\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t)&0\\ 0&(\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t))^{\ast}\end{array}\right]$ (51)
    and $\widetilde{U}^{\ast}_{\rho}(t)\Delta\widetilde{U}_{\rho}(t)=t^{1-1/\rho}\Delta_{\rho}$, where $\widetilde{\Theta}_{\rho}(t)=\Theta_{\rho}(S_{\rho})+O(t^{1/\rho})$ and $\Delta_{\rho}=\left[\begin{array}[]{cc}0&I_{\rho s_{\rho}}\\ I_{\rho s_{\rho}}&0\end{array}\right]$.
Proof. Part (a) is directly from Lemma 1.
For part (b), by applying Lemma 1 to $\widetilde{C}_{1}(t)$ and using (34), one has
$(C+tD)\widetilde{V}_{\rho}(t)=\widetilde{V}_{\rho}(t)(\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t)),$
where $\widetilde{\Theta}_{\rho}(t)=\Theta_{\rho}(S_{\rho})+O(t^{1/\rho})$ and
$\widetilde{V}_{\rho}(t)=UZ(t)\left[\begin{array}[]{c}\widetilde{Z}_{1}(t)\left[\begin{array}[]{c}X_{\rho}(t)\\ 0\end{array}\right]\\ 0\end{array}\right]$ (52)
with $X_{\rho}(t)$ given in (5). Since $Z(t)=I_{2n}+O(t)$ and $\widetilde{Z}_{1}(t)=I_{2p}+O(t)$,
$\widetilde{V}_{\rho}(t)=U\left[\begin{array}[]{c}X_{\rho}(t)\\ 0\\ 0\end{array}\right]+O(t)=VX_{\rho}(t)+O(t).$
The formula for $\widetilde{V}_{\rho}(t)$ is derived by using the block structures of $X_{\rho}(t)$ and $V$.
For the block $\widetilde{C}_{1}^{\ast}(t)$, we apply Lemma 1 to $P^{\ast}\widetilde{C}_{1}^{\ast}(t)P$. Based on (47), the matrix corresponding to $W_{k}$ is $\widetilde{W}_{k}=W_{k}^{\ast}$ defined in (48). Hence, the matrix corresponding to $S_{\rho}$ is $S_{\rho}^{\ast},$ and we have
$(C+tD)\widehat{V}_{\rho}^{(c)}(t)=\widehat{V}_{\rho}^{(c)}(t)(P_{\rho}^{\ast}(\bar{\lambda}I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}^{(c)}(t))P_{\rho}),$
where $\widetilde{\Theta}_{\rho}^{(c)}(t)=\Theta_{\rho}(S_{\rho}^{\ast})+O(t^{1/\rho})$,
$\widehat{V}_{\rho}^{(c)}(t)=UZ(t)\left[\begin{array}[]{c}\widetilde{Z}_{1}(t)\left[\begin{array}[]{c}0\\ P^{\ast}X^{(c)}_{\rho}(t)P_{\rho}\end{array}\right]\\ 0\end{array}\right]=V_{c}P^{\ast}X^{(c)}_{\rho}(t)P_{\rho}+O(t),$ (53)
$P$ and $P_{\rho}$ are defined in (46), and $X^{(c)}_{\rho}(t)$ has the same pattern as $X_{\rho}(t)$ with $G_{\rho}$ replaced by $G_{\rho}^{(c)}$. Partition $P^{\ast}X^{(c)}_{\rho}(t)P_{\rho}=\left[\begin{array}[]{ccc}(X_{1\rho}^{(c)})^{T}(t)&\ldots&(X_{m\rho}^{(c)})^{T}(t)\end{array}\right]^{T}$. One has
$X^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cc}0&0\\ 0&I_{s_{\rho}}\end{array}\right]+\mathop{\rm diag}\nolimits(t^{1-1/\rho}I_{s_{\rho}},\ldots,t^{1/\rho}I_{s_{\rho}},0_{s_{\rho}})&k=\rho\\ \left[\begin{array}[]{cc}0&0\\ 0&G_{k\rho}^{(c)}\end{array}\right]+O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho.\end{array}\right$.
Using the $\Delta$-Hermitian property of $C+tD$, one has
$(\Delta\widetilde{V}_{\rho}(t))^{\ast}(C+tD)$ $=$ $(\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t))^{\ast}(\Delta\widetilde{V}_{\rho}(t))^{\ast}$
$(\Delta\widehat{V}_{\rho}^{(c)}(t))^{\ast}(C+tD)$ $=$ $(P_{\rho}^{\ast}(\bar{\lambda}I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}^{(c)}(t))P_{\rho})^{\ast}(\Delta\widehat{V}_{\rho}^{(c)}(t))^{\ast}.$
This indicates that $\Delta\widetilde{V}_{\rho}(t)$ is a basis matrix of the left invariant subspace of all the eigenvalues of $C+tD$ that are perturbed from $\bar{\lambda}$ of order $O(t^{1/\rho})$. Similarly, $\Delta\widehat{V}_{\rho}^{(c)}(t)$ is a basis matrix of the left invariant subspace of all the eigenvalues of $C+tD$ that are perturbed from $\lambda$ of order $O(t^{1/\rho})$. Therefore,
$(\Delta\widetilde{V}_{\rho}(t))^{\ast}\widetilde{V}_{\rho}(t)=(\Delta\widehat{V}_{\rho}^{(c)}(t))^{\ast}\widehat{V}_{\rho}^{(c)}(t)=0,$
and $(\Delta\widetilde{V}_{\rho}(t))^{\ast}\widehat{V}_{\rho}^{(c)}(t)$ is invertible. Using (25), (52), and (53),
$\widetilde{V}_{\rho}^{\ast}(t)\Delta\widehat{V}_{\rho}^{(c)}(t)$ $=$ $\left(UZ(t)\left[\begin{array}[]{c}\widetilde{Z}_{1}(t)\left[\begin{array}[]{c}X_{\rho}(t)\\ 0\end{array}\right]\end{array}\right]\right)^{\ast}\Delta\left(UZ(t)\left[\begin{array}[]{c}\widetilde{Z}_{1}(t)\left[\begin{array}[]{c}0\\ P^{\ast}X_{\rho}^{(c)}(t)P_{\rho}\end{array}\right]\\ 0\end{array}\right]\right)$ (61)
$=$ $\left(\widetilde{Z}_{1}(t)\left[\begin{array}[]{c}X_{\rho}(t)\\ 0\end{array}\right]\right)^{\ast}\Delta_{T}\left(\widetilde{Z}_{1}(t)\left[\begin{array}[]{c}0\\ P^{\ast}X_{\rho}^{(c)}(t)P_{\rho}\end{array}\right]\right)$ (66)
$=$ $\left[\begin{array}[]{c}X_{\rho}(t)\\ 0\end{array}\right]^{\ast}\Delta_{T}\left[\begin{array}[]{c}0\\ P^{\ast}X_{\rho}^{(c)}(t)P_{\rho}\end{array}\right]=X_{\rho}^{\ast}(t)P^{\ast}X_{\rho}^{(c)}(t)P_{\rho}$
$=$ $\sum_{k=1}^{m}X_{k\rho}^{\ast}(t)X_{k\rho}^{(c)}(t)=t^{1-1/\rho}I_{\rho s_{\rho}}+O(t)=t^{1-1/\rho}R_{\rho}(t),$
where $R_{\rho}(t)=I_{\rho s_{\rho}}+O(t^{1/\rho})$. Let
$\widehat{X}_{\rho}^{(c)}(t)=P^{\ast}X_{\rho}^{(c)}(t)P_{\rho}R_{\rho}^{-1}(t),\qquad\widetilde{V}_{\rho}^{(c)}(t)=\widehat{V}_{\rho}^{(c)}(t)R_{\rho}^{-1}(t)=V_{c}\widehat{X}_{\rho}^{(c)}(t)+O(t),$ (72)
and $\widetilde{U}_{\rho}(t)=\left[\begin{array}[]{cc}\widetilde{V}_{\rho}(t)&\widetilde{V}_{\rho}^{(c)}(t)\end{array}\right]$. One has
$\widetilde{U}_{\rho}^{\ast}(t)\Delta\widetilde{U}_{\rho}(t)=\left[\begin{array}[]{cc}0&\widetilde{V}_{\rho}^{\ast}(t)\Delta\widetilde{V}_{\rho}^{(c)}(t)\\ (\widetilde{V}_{\rho}^{\ast}(t)\Delta\widetilde{V}_{\rho}^{(c)}(t))^{\ast}&0\end{array}\right]=t^{1-1/\rho}\Delta_{\rho}.$
Since $R_{\rho}^{-1}(t)=I_{\rho s_{\rho}}+O(t^{1/\rho})$, partition $\widehat{X}_{\rho}^{(c)}(t)=\left[\begin{array}[]{ccc}(\widehat{X}_{1\rho}^{(c)})^{T}(t)&\ldots&(\widehat{X}_{m\rho}^{(c)})^{T}(t)\end{array}\right]^{T}$ conformably with $P^{\ast}X_{\rho}^{(c)}(t)P_{\rho}$, one has
$\widehat{X}^{(c)}_{k\rho}(t)=X^{(c)}_{k\rho}(t)R_{\rho}^{-1}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cc}0&0\\ 0&I_{\rho}\end{array}\right]+\mathop{\rm diag}\nolimits(t^{1-1/\rho}I_{s_{\rho}},\ldots,t^{1/\rho}I_{s_{\rho}},0_{s_{\rho}})&\\ \qquad+O(\left[\begin{array}[]{cccc}t&t^{1-1/\rho}&\ldots&t^{1/\rho}\end{array}\right]^{T}),&k=\rho\\ \left[\begin{array}[]{cc}0&0\\ 0&G_{k\rho}^{(c)}\end{array}\right]+O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho,\end{array}\right$.
with which the expression of $\widetilde{V}_{\rho}^{(c)}(t)$ can be derivd.
Finally,
$(C+tD)\widetilde{U}_{\rho}(t)=\widetilde{U}_{\rho}(t)\left[\begin{array}[]{cc}\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t)&0\\ 0&R_{\rho}(t)P_{\rho}^{\ast}(\bar{\lambda}I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}^{(c)}(t))P_{\rho}R_{\rho}^{-1}(t)\end{array}\right].$
Because
$\widetilde{U}_{\rho}^{\ast}(t)\Delta(C+tD)\widetilde{U}_{\rho}(t)$
$\qquad=\widetilde{U}_{\rho}^{\ast}(t)\Delta\widetilde{U}_{\rho}(t)\left[\begin{array}[]{cc}\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t)&0\\ 0&R_{\rho}(t)P_{\rho}^{\ast}(\bar{\lambda}I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}^{(c)}(t))P_{\rho}R_{\rho}^{-1}(t)\end{array}\right]$
$\qquad=t^{1-1/\rho}\Delta_{\rho}\left[\begin{array}[]{cc}\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t)&0\\ 0&R_{\rho}(t)P_{\rho}^{\ast}(\bar{\lambda}I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}^{(c)}(t))P_{\rho}R_{\rho}^{-1}(t)\end{array}\right]$
is Hermitian, one has
$R_{\rho}(t)P_{\rho}^{\ast}(\bar{\lambda}I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}^{(c)}(t))P_{\rho}R_{\rho}^{-1}(t)=(\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t))^{\ast}.$
The next result is about the perturbations of the invariant subspaces corresponding to subsets of $\Lambda(\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t))$ and $\Lambda((\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t))^{\ast})$.
Suppose $F,F^{(c)}\in{\mathbb{C}}^{\rho s_{\rho}\times\ell}$ are full column rank matrices satisfying
$\Theta_{\rho}(S_{\rho})F=F\Omega,\quad\Theta_{\rho}^{\ast}(S_{\rho})F^{(c)}=F^{(c)}\Omega^{\ast},\quad F^{\ast}F^{(c)}=I_{\ell},$ (75)
for some $\Omega\in{\mathbb{C}}^{\ell\times\ell}$ that satisfies
$\Lambda(\Omega)\subseteq\Lambda(\Theta_{\rho}(S_{\rho})),\qquad\Lambda(\Omega)\cap(\Lambda(\Theta_{\rho}(S_{\rho}))\backslash\Lambda(\Omega))=\emptyset.$ (76)
Due to the block structure of $\Theta_{\rho}(S_{\rho})$, $F$ and $F^{(c)}$ have the block forms
$F=\left[\begin{array}[]{cccc}Q^{T}&(Q\Omega)^{T}&\ldots&(Q\Omega^{\rho-1})^{T}\end{array}\right]^{T},\quad F^{(c)}=\left[\begin{array}[]{cccc}\Omega^{\rho-1}(Q^{(c)})^{\ast}&\ldots&\Omega(Q^{(c)})^{\ast}&(Q^{(c)})^{\ast}\end{array}\right]^{\ast},$
for some matrices $Q,Q^{(c)}\in{\mathbb{C}}^{s_{\rho}\times\ell}$ satisfying
$S_{\rho}Q=Q\Omega^{\rho},\quad S_{\rho}^{\ast}Q^{(c)}=Q^{(c)}(\Omega^{\ast})^{\rho}.$
Theorem 6
Under the conditions of Theorem 5, suppose $F$ and $F^{(c)}$ are the full column rank matrices satisfying (75), and $\Omega$ satisfies (76). For $t$ sufficiently small, there exist matrices $\widetilde{F}(t)=F+O(t^{1/\rho})$, $\widetilde{F}^{(c)}(t)=F^{(c)}+O(t^{1/\rho})$, and $\widetilde{\Omega}(t)=\Omega+O(t^{1/\rho})$ satisfying
$\widetilde{\Theta}_{\rho}(t)\widetilde{F}(t)=\widetilde{F}(t)\widetilde{\Omega}(t),\quad\widetilde{\Theta}_{\rho}^{\ast}(t)\widetilde{F}^{(c)}(t)=\widetilde{F}^{(c)}(t)\widetilde{\Omega}^{\ast}(t),\quad\widetilde{F}^{\ast}(t)\widetilde{F}^{(c)}(t)=I_{\ell},$
such that the matrix $\widetilde{U}_{\Omega}(t)=\left[\begin{array}[]{cc}L_{\rho}(t)&L_{\rho}^{(c)}(t)\end{array}\right]$ satisfies
$(C+tD)\widetilde{U}_{\Omega}(t)=\widetilde{U}_{\Omega}(t)\left[\begin{array}[]{cc}\lambda I_{\ell}+t^{1/\rho}\widetilde{\Omega}(t)&0\\ 0&(\lambda I_{\ell}+t^{1/\rho}\widetilde{\Omega}(t))^{\ast}\end{array}\right],\qquad\widetilde{U}^{\ast}_{\Omega}(t)\Delta\widetilde{U}_{\Omega}(t)=t^{1-1/\rho}\Delta_{\Omega},$
where $\Delta_{\Omega}=\left[\begin{array}[]{cc}0&I_{\ell}\\ I_{\ell}&0\end{array}\right],$ and
$L_{\rho}(t)$ $=$ $\widetilde{V}_{\rho}(t)\widetilde{F}(t)=E_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ G_{\rho}\end{array}\right]Q+\sum_{k=1}^{m}V_{k}L_{k\rho}(t)+O(t),$
$L_{\rho}^{(c)}$ $=$ $\widetilde{V}^{(c)}_{\rho}(t)\widetilde{F}^{(c)}(t)=E_{\rho}^{(c)}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ G_{\rho}^{(c)}\end{array}\right]Q^{(c)}+\sum_{k=1}^{m}V_{k}^{(c)}L_{k\rho}^{(c)}(t)+O(t)$
with
$L_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
and
$L^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\Omega)^{\rho-1}(Q^{(c)})^{\ast}&\ldots&(t^{1/\rho}\Omega)(Q^{(c)})^{\ast}&0\end{array}\right]^{\ast}&\\ \qquad\qquad\qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho.\end{array}\right$.
Proof. Because
$\widetilde{\Theta}_{\rho}(t)=\Theta_{\rho}(S_{\rho})+O(t^{1/\rho}),$
following the standard perturbation theory ([26]), for $t$ sufficiently small, there are
$\widetilde{F}(t)=F+O(t^{1/\rho}),\quad\widehat{F}^{(c)}(t)=F^{(c)}+O(t^{1/\rho})$
such that
$\widetilde{\Theta}_{\rho}(t)\widetilde{F}(t)=\widetilde{F}(t)\widetilde{\Omega}(t),\quad\widetilde{\Theta}_{\rho}^{\ast}(t))\widehat{F}^{(c)}(t)=\widehat{F}^{(c)}(t)\widehat{\Omega}^{\ast}(t),$
with $\widetilde{\Omega}(t)=\Omega+O(t^{1/\rho}),$ $\widehat{\Omega}(t)=\Omega+O(t^{1/\rho}).$ Let
$\widehat{R}(t)=\widetilde{F}^{\ast}(t)\widehat{F}^{(c)}(t)=F^{\ast}F^{(c)}+O(t^{1/\rho})=I_{\ell}+O(t^{1/\rho}).$
Define
$\widetilde{F}^{(c)}(t)=\widehat{F}^{(c)}(t)\widehat{R}^{-1}(t).$
Then
$\widetilde{F}^{\ast}(t)\widetilde{F}^{(c)}(t)=I_{\ell},\quad\widetilde{F}^{(c)}(t)=F^{(c)}+O(t^{1/\rho}).$
Let $L_{\rho}(t)=\widetilde{V}_{\rho}(t)\widetilde{F}(t)$, $L_{\rho}^{(c)}(t)=\widetilde{V}_{\rho}^{(c)}(t)\widetilde{F}^{(c)}(t)$, and $\widetilde{U}_{\Omega}(t)=\left[\begin{array}[]{cc}L_{\rho}(t)&L_{\rho}^{(c)}(t)\end{array}\right]$. The results can be proved by using (51).
When $\Omega$ is a scalar, we have the following result.
Corollary 7
Under the conditions of Theorem 5, suppose $\omega$ is a simple eigenvalue of $\Theta_{\rho}(S_{\rho})$. The vectors
$f=\left[\begin{array}[]{cccc}q^{T}&\omega q^{T}&\ldots&\omega^{\rho-1}q^{T}\end{array}\right]^{T},\quad f^{(c)}=\left[\begin{array}[]{cccc}\omega^{\rho-1}(q^{(c)})^{\ast}&\ldots&\omega(q^{(c)})^{\ast}&(q^{(c)})^{\ast}\end{array}\right]^{\ast},$
satisfy
$\Theta_{\rho}(S_{\rho})f=\omega f,\quad\Theta_{\rho}^{\ast}(S_{\rho})f^{(c)}=\bar{\omega}f^{(c)};\quad S_{\rho}q=\omega^{\rho}q,\quad S_{\rho}^{\ast}q^{(c)}=\bar{\omega}^{\rho}q^{(c)},$
and
$f^{\ast}f^{(c)}=\rho{\bar{\omega}}^{\rho-1}q^{\ast}q^{(c)}=1.$
For $t$ sufficiently small, there are vectors $\widetilde{f}(t)=f+O(t^{1/\rho})$, $\widetilde{f}^{(c)}(t)=f^{(c)}+O(t^{1/\rho})$, and a scalar $\widetilde{\omega}(t)=\omega+O(t^{1/\rho})$ satisfying
$\widetilde{\Theta}_{\rho}(t)\widetilde{f}(t)=\widetilde{\omega}(t)\widetilde{f}(t),\quad\widetilde{\Theta}_{\rho}^{\ast}(t)\widetilde{f}^{(c)}=\overline{\widetilde{\omega}(t)}\widetilde{f}^{(c)}(t),\quad\widetilde{f}^{\ast}(t)\widetilde{f}^{(c)}(t)=1,$
such that matrix $\widetilde{U}_{\omega}(t)=\left[\begin{array}[]{cc}l_{\rho}(t)&l_{\rho}^{(c)}(t)\end{array}\right]$ satisfies
$(C+tD)\widetilde{U}_{\omega}(t)=\widetilde{U}_{\omega}(t)\left[\begin{array}[]{cc}\lambda+t^{1/\rho}\widetilde{\omega}(t)&0\\ 0&\overline{\lambda+t^{1/\rho}\widetilde{\omega}(t)}\end{array}\right],\quad\widetilde{U}_{\omega}^{\ast}(t)\Delta\widetilde{U}_{\omega}(t)=t^{1-1/\rho}\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right],$
where
$l_{\rho}(t)$ $=$ $\widetilde{V}_{\rho}(t)\widetilde{f}(t)=E_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ G_{\rho}\end{array}\right]q+\sum_{k=1}^{m}V_{k}l_{k\rho}(t)+O(t),$
$l_{\rho}^{(c)}(t)$ $=$ $\widetilde{V}_{\rho}^{(c)}(t)\widetilde{f}^{(c)}(t)=E_{\rho}^{(c)}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ G_{\rho}^{(c)}\end{array}\right]q^{(c)}+\sum_{k=1}^{m}V_{k}^{(c)}l_{k\rho}^{(c)}(t)+O(t)$
with
$l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\omega q^{T}&\ldots&(t^{1/\rho}\omega)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho,\end{array}\right$.
$l^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\omega)^{\rho-1}(q^{(c)})^{\ast}&\ldots&t^{1/\rho}\omega(q^{(c)})^{\ast}&0\end{array}\right]^{\ast}+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho.\end{array}\right$.
Proof. It is a special case of Theorem 6 with $\ell=1$.
Remark 2
Based on (34), the matrix $U\left[\begin{array}[]{c}I\\ Y(t)\end{array}\right]Z_{1}(t)\widetilde{Z}_{1}(t)$ is a basis matrix of the invariant subspace of $C+tD$ corresponding to the eigenvalues of $\mathop{\rm diag}\nolimits(\widetilde{C}_{1}(t),\widetilde{C}_{1}^{\ast}(t))$ and it satisfies
$\left(U\left[\begin{array}[]{c}I\\ Y(t)\end{array}\right]Z_{1}(t)\widetilde{Z}_{1}(t)\right)^{\ast}\Delta\left(U\left[\begin{array}[]{c}I\\ Y(t)\end{array}\right]Z_{1}(t)\widetilde{Z}_{1}(t)\right)=\Delta_{T}.$
If we apply Theorem 5 to every $\rho$ from $1$ to $m$, we are able to assemble another basis matrix $\widetilde{U}(t)=\left[\begin{array}[]{ccc}\widetilde{U}_{1}(t),\ldots,\widetilde{U}_{m}(t)\end{array}\right]$ corresponding to the eigenvalues of $\mathop{\rm diag}\nolimits(\widetilde{C}_{1}(t),\widetilde{C}_{1}^{\ast}(t))$ as well. However, these two basis matrices are different. The former basis matrix corresponds to $\widetilde{C}_{1}(t)$ and $\widetilde{C}_{1}^{\ast}(t)$ that are perturbed from $\lambda I_{p}+N$ and $(\lambda I_{p}+N)^{\ast}$ with an order $O(t)$ perturbation. On the other hand, $\widetilde{U}(t)$ corresponds to a block diagonal matrix with diagonal blocks perturbed from $\lambda I_{\rho s_{\rho}}$ and $\bar{\lambda}I_{\rho s_{\rho}}$ with perturbations of orders $O(t^{1/\rho})$ for $\rho=1,\ldots,m$. Meanwhile $\widetilde{U}_{\rho}(t)=O(1)$, for $1\leq\rho\leq m$, and
$\widetilde{U}^{\ast}(t)\Delta\widetilde{U}(t)=\mathop{\rm diag}\nolimits(\Delta_{1},t^{1/2}\Delta_{2},\ldots,t^{1-1/m}\Delta_{m}).$
The former basis matrix simply considers all the eigenvalues perturbed from $\lambda$ and $\bar{\lambda}$ as a single set. In contrast, $\widetilde{U}(t)$ provides information about every individual perturbed eigenvalue, which is shown in Theorem 6 and Corollary 7. In fact, the proofs of Theorem 6 and Corollary 7 reveal that the perturbation formulas for smaller invariant subspaces and eigenvectors can be derived from $\widetilde{U}(t)$ with transformations that are essentially constant.
Remark 3
In Corollary 7, $l_{\rho}(t)$ and $l_{\rho}^{(c)}(t)$ are right and left eigenvectors of $C+tD$ corresponding to the eigenvalue $\lambda+t^{1/\rho}\widetilde{w}(t)$. Note that $l_{\rho}(t),l_{\rho}^{(c)}(t)=O(1)$ and $l_{\rho}^{\ast}(t)l_{\rho}^{(c)}(t)=t^{1-1/\rho}$. The latter indicates that the spectral condition numbers of $C+tD$ corresponding to the eigenvalue $\lambda+t^{1/\rho}\widetilde{\omega}(t)$ is $O(t^{-(1-1/\rho)})$. If we consider $C$ as the matrix perturbed from $C+tD$, since the spectral condition number is $O(t^{-(1-1/\rho)})$, the eigenvalue should change with an order $O(t^{1/\rho})$. In this sense, Corollary 7 generalizes the standard first order perturbation results for simple eigenvalues ([13, Sec.7.2.2]).
3.2 Real eigenvalue case
In this case, we consider (19), where $\widetilde{C}_{T}(t)=C_{T}+t\widetilde{D}_{11}(t)$,
$C_{T}=\alpha I_{p}+N,\quad\widetilde{D}_{11}(t)=D_{11}+O(t)=\Delta_{T}U_{T}^{\ast}\Delta DU_{T}+O(t),$
$\alpha\in{\mathbb{R}}$, and $\Delta_{T}$ is given in (9). Recall that $C_{T}$, $D_{11}$, and $\widetilde{D}_{11}(t)$ are $\Delta_{T}$-Hermitian. Partition
$U_{T}=\left[\begin{array}[]{cccc}U_{1}&U_{2}&\ldots&U_{m}\end{array}\right]$ (81)
and
$U_{j}=\left[\begin{array}[]{cccc}U_{j1}&U_{j2}&\ldots&U_{jj}\end{array}\right],\quad j=1,\ldots,m,$
conformably with the block columns of $N$ and $N_{j}$, respectively. Define
$E_{k}=\left[\begin{array}[]{cccc}U_{k1}&U_{k+1,1}&\ldots&U_{m1}\end{array}\right],\quad k=1,\ldots,m.$
Then
$D_{11}=\Delta_{T}U_{T}^{\ast}\Delta DU_{T}=\Delta_{T}\left[\begin{array}[]{cccc}U_{1}^{\ast}\Delta DU_{1}&U_{1}^{\ast}\Delta DU_{2}&\ldots&U_{1}^{\ast}\Delta DU_{m}\\ U_{2}^{\ast}\Delta DU_{1}&U_{2}^{\ast}\Delta DU_{2}&\ldots&U_{2}^{\ast}\Delta DU_{m}\\ \vdots&\vdots&\ddots&\vdots\\ U_{m}^{\ast}\Delta DU_{1}&U_{m}^{\ast}\Delta DU_{2}&\ldots&U_{m}^{\ast}\Delta DU_{m}\end{array}\right].$
If we consider $B(t)$ and $B$ in Lemma 1 as $\widetilde{D}_{11}(t)$ and $D_{11}$, one has
$W_{k}=\left[\begin{array}[]{ccc}\Sigma_{k}&&\\ &\ddots&\\ &&\Sigma_{m}\end{array}\right]\widehat{W}_{k},\quad\widehat{W}_{k}=E_{k}^{\ast}\Delta DE_{k}=\left[\begin{array}[]{cc}U_{k1}^{\ast}\Delta DU_{k1}&U_{k1}^{\ast}\Delta DE_{k+1}\\ E_{k+1}^{\ast}\Delta DU_{k1}&\widehat{W}_{k+1}\end{array}\right]$ (82)
for $k=1,2,\ldots,m-1$, and $\widehat{W}_{m}=U_{m1}^{\ast}\Delta DU_{m1}$. Note $\widehat{W}_{k}$ is Hermitian for any $k$.
Theorem 8
Assume both $C$ and $D$ are $\Delta$-Hermitian matrices and $C+tD$ has the decomposition (19) when $t$ is sufficiently small, with $C_{T}=\alpha I_{p}+N$, $\alpha\in{\mathbb{R}}$, and $\Delta_{T}$ defined in (9). Let $W_{k}$ be defined as in (82) for $k\in\{1,\ldots,m\}$. For a fixed $\rho\in\{1,\ldots,m\}$ assume $W_{\rho+1}$ is invertible if $\rho<m$. Define
$G_{\rho}=-\widehat{W}_{\rho+1}^{-1}E_{\rho+1}^{\ast}\Delta DU_{\rho 1},$
and let $\gamma_{1}^{(\rho)},\ldots,\gamma_{s_{\rho}}^{(\rho)}$ be the $s_{\rho}$ eigenvalues of the matrix
$S_{\rho}=\Sigma_{\rho}\widehat{S}_{\rho},\quad\mbox{where}\quad\widehat{S}_{\rho}=\widehat{S}_{\rho}^{\ast}=\left\{\begin{array}[]{ll}U_{\rho 1}^{\ast}\Delta D(U_{\rho 1}+E_{\rho+1}G_{\rho})&\rho<m\\ U_{m1}^{\ast}\Delta DU_{m1}&\rho=m.\end{array}\right$.
  1. (a)
[…]
    $\lambda_{ij}^{(\rho)}(t)=\alpha+t^{1/\rho}\mu_{ij}^{(\rho)}+o(t^{1/\rho}),\qquad i=1,\ldots,s_{\rho},\quad j=1,\ldots,\rho.$
  2. (b)
    Suppose $W_{1},\ldots,W_{m}$ are all invertible. Let $\Theta_{\rho}(S_{\rho})$ be defined as (4) with $S_{\rho}$ given above. There is a matrix
    $\widetilde{U}_{\rho}(t)=E_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}&0\\ G_{\rho}&0\end{array}\right]+\sum_{k=1}^{m}U_{k}\widetilde{X}_{k\rho}(t)+O(t),\\$
    satisfying
    $(C+tD)\widetilde{U}_{\rho}(t)=\widetilde{U}_{\rho}(t)(\alpha I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t)),\qquad\widetilde{\Theta}_{\rho}(t)=\Theta_{\rho}(S_{\rho})+O(t^{1/\rho})$
    and $\widetilde{U}_{\rho}^{\ast}(t)\Delta\widetilde{U}_{\rho}(t)=t^{1-1/\rho}\Gamma_{\rho}$, where
    $\widetilde{X}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \mathop{\rm diag}\nolimits(0_{s_{\rho}},t^{1/\rho}I_{s_{\rho}},\ldots,t^{1-1/\rho}I_{s_{\rho}})+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho,\end{array}\right$.
    $U_{1},\ldots,U_{m}$ are the block columns of $U_{T}$ given in (81), and $\Gamma_{\rho}$ is defined in Proposition 3.
Proof. From (19) one has
$(C+tD)U\left[\begin{array}[]{c}I\\ Y(t)\end{array}\right]Z_{1}(t)=U\left[\begin{array}[]{c}I\\ Y(t)\end{array}\right]Z_{1}(t)\widetilde{C}_{T}(t),$
$\left(U\left[\begin{array}[]{c}I\\ Y(t)\end{array}\right]Z_{1}(t)\right)^{\ast}\Delta\left(U\left[\begin{array}[]{c}I\\ Y(t)\end{array}\right]Z_{1}(t)\right)=\Delta_{T},$
where $Y(t)=O(t)$, $Z_{1}(t)=I_{r}+O(t^{2})$, and $\widetilde{C}_{T}(t)=C_{T}+t\widetilde{D}_{11}(t)$, $\widetilde{D}_{11}(t)=D_{11}+O(t)$, $D_{11}=\Delta_{T}U_{T}^{\ast}\Delta DU_{T}$ satisfying
$(\Delta_{T}\widetilde{C}_{T}(t))^{\ast}=\Delta_{T}\widetilde{C}_{T}(t),\quad(\Delta_{T}D_{11})^{\ast}=\Delta_{T}D_{11}.$
Applying Lemma 1(a) to $\widetilde{C}_{T}(t)$ with $B(t)=\widetilde{D}_{11}(t)$ and $B=D_{11}$, we have part (a).
For part (b), following Lemma 1(b) one has
$(C+tD)\widehat{U}_{\rho}(t)=\widehat{U}_{\rho}(t)(\alpha I_{p}+t^{1/\rho}\widehat{\Theta}_{\rho}(t)),$
where $\widehat{\Theta}_{\rho}(t)=\Theta_{\rho}(S_{\rho})+O(t^{1/\rho})$,
$\widehat{U}_{\rho}(t)=U\left[\begin{array}[]{c}I\\ Y(t)\end{array}\right]Z_{1}(t)X_{\rho}(t),$
and $X_{\rho}(t)$ is defined as in (5). Using (19),
$\widehat{U}_{\rho}^{\ast}(t)\Delta\widehat{U}_{\rho}(t)$ $=$ $X_{\rho}^{\ast}(t)\Delta_{T}X_{\rho}(t)=X_{\rho\rho}^{\ast}(t)\Gamma_{\rho}X_{\rho\rho}(t)+\sum_{k\neq\rho}X_{k\rho}^{\ast}(t)\Gamma_{k}X_{k\rho}(t)$
$=$ $t^{1-1/\rho}\Gamma_{\rho}+O(t)=t^{1-1/\rho}\widetilde{R}_{\rho}(t),$
where $\widetilde{R}_{\rho}(t)=\Gamma_{\rho}+O(t^{1/\rho})$. Note $\widetilde{R}_{\rho}(t)$ is Hermitian. As did in the beginning of the section for (14), when $t$ is sufficiently small, one has $Z_{\rho}(t)=I+O(t^{1/\rho})$ such that
$Z_{\rho}^{\ast}(t)\widetilde{R}_{\rho}(t)Z_{\rho}(t)=\Gamma_{\rho}.$
Let
$\widetilde{U}_{\rho}(t)=\widehat{U}_{\rho}(t)Z_{\rho}(t)=U_{T}X_{\rho}(t)Z_{\rho}(t)+O(t),\quad\widetilde{\Theta}_{\rho}(t)=Z_{\rho}^{-1}(t)\widehat{\Theta}_{\rho}(t)Z_{\rho}(t)=\Theta_{\rho}(S_{\rho})+O(t^{1/\rho}).$
The results follow.
Remark 4
From Theorem 8(b), $\widetilde{\Theta}_{\rho}(t)$ is $\Gamma_{\rho}$-Hermitian. Therefore, the eigenvalues of $\widetilde{\Theta}_{\rho}(t)$, or equivalently the eigenvalues $\lambda_{ij}^{(\rho)}(t)$, are either in conjugate pairs or real. Since in this case $S_{\rho}$ is $\Sigma_{\rho}$-Hermitian, $\Theta_{\rho}(S_{\rho})$ is $\Gamma_{\rho}$-Hermitian as well. When $t$ is sufficiently small, the real eigenvalues of $\widetilde{\Theta}_{\rho}(t)$ are perturbed from those real $\mu_{ij}^{(\rho)}$’s, which correspond to the real $\gamma_{i}^{(\rho)}$’s when $\rho$ is odd or real positive $\gamma_{i}^{(\rho)}$’s when $\rho$ is even. However, because of the relation $\widetilde{\Theta}_{\rho}(t)=\Theta_{\rho}(S_{\rho})+O(t^{1/\rho})$, not all real $\mu_{ij}^{(\rho)}$’s correspond to the real eigenvalues of $\widetilde{\Theta}_{\rho}(t)$. The $O(t^{1/\rho})$ term could introduce nonzero imaginary parts.
For the perturbations of invariant subspaces with smaller dimensions, because $\Theta_{\rho}(S_{\rho})$ and $\widetilde{\Theta}_{\rho}(t)$ are $\Gamma_{\rho}$-Hermitian matrices, we will consider the situation with a full column rank matrix $F\in{\mathbb{C}}^{\rho s_{\rho}\times\ell}$ such that
$\Theta_{\rho}(S_{\rho})F=F\Omega,\quad F^{\ast}\Gamma_{\rho}F=\Delta_{\Omega},$ (83)
where $\Omega\in{\mathbb{C}}^{\ell\times\ell}$ with the eigenvalues either real or in complex conjugate pairs, $\Lambda(\Omega)\cap(\Lambda(\Theta_{\rho}(S_{\rho}))\backslash\Lambda(\Omega))=\emptyset$, and $\Delta_{\Omega}$ is Hermitian and block diagonal as $\widehat{\Delta}$ defined in Proposition 3. Due to the block structure of $\Theta_{\rho}(S_{\rho})$, it still has
$F=\left[\begin{array}[]{cccc}Q^{T}&(Q\Omega)^{T}&\ldots&(Q\Omega^{\rho-1})^{T}\end{array}\right]^{T}$ (84)
for some matrix $Q\in{\mathbb{C}}^{s_{\rho}\times\ell}$ satisfying $S_{\rho}Q=Q\Omega^{\rho}$. Using the same method as we did in the beginning of the section, we can show the existence of
$\widetilde{F}(t)=F+O(t^{1/\rho}),\qquad\widetilde{\Omega}(t)=\Omega+O(t^{1/\rho})$ (85)
satisfying
$\widetilde{\Theta}_{\rho}(t)\widetilde{F}(t)=\widetilde{F}(t)\widetilde{\Omega}(t),\quad\widetilde{F}^{\ast}(t)\Gamma_{\rho}\widetilde{F}(t)=\Delta_{\Omega},$ (86)
Theorem 9
Under the conditions of Theorem 8, suppose $F$ is a full column rank matrix of the form (84) that satisfies (83). Also, suppose $\Lambda(\Omega)\subseteq\Lambda(\Theta_{\rho}(S_{\rho}))$ with the eigenvalues in $\Lambda(\Omega)$ being either real or in complex conjugate pairs and $\Lambda(\Omega)\cap(\Lambda(\Theta_{\rho}(S_{\rho}))\backslash\Lambda(\Omega))=\emptyset$. Then for $t$ sufficiently small, there exist matrices $\widetilde{F}(t)$ and $\widetilde{\Omega}(t)$ satisfying (85) and (86), such that the matrix
$\widetilde{U}_{\Omega}(t)=\widetilde{U}_{\rho}(t)\widetilde{F}(t)=E_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ G_{\rho}\end{array}\right]Q+\sum_{k=1}^{m}U_{k}L_{k\rho}(t)+O(t)$
with
$L_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
satisfies
$(C+tD)\widetilde{U}_{\Omega}(t)=\widetilde{U}_{\Omega}(t)(\alpha I_{\ell}+t^{1/\rho}\widetilde{\Omega}(t)),\qquad\widetilde{U}^{\ast}_{\Omega}(t)\Delta\widetilde{U}_{\Omega}(t)=t^{1-1/\rho}\Delta_{\Omega}.$
Proof. The results can be derived by setting $\widetilde{U}_{\Omega}(t)=\widetilde{U}_{\rho}(t)\widetilde{F}(t),$ where $\widetilde{U}_{\rho}(t)$ is given in Theorem 8, and using (85) and (86).
The following results are the special cases when either $\Omega$ is a simple real eigenvalue of $\Theta_{\rho}(S_{\rho})$, or $\Omega=\left[\begin{array}[]{cc}\mu&0\\ 0&\bar{\mu}\end{array}\right]$, where $\mu,\bar{\mu}$ are nonreal simple eigenvalues of $\Theta_{\rho}(S_{\rho})$.
Corollary 10
  1. (a)
    Suppose $\omega$ is a simple real eigenvalue of $\Theta_{\rho}(S_{\rho})$. There is a vector
    $f=\left[\begin{array}[]{cccc}q^{T}&\omega q^{T}&\ldots&\omega^{\rho-1}q^{T}\end{array}\right]^{T}$ (87)
    that satisfies
    $\Theta_{\rho}(S_{\rho})f=\omega f,\quad S_{\rho}q=\omega^{\rho}q,\quad f^{\ast}\Gamma_{\rho}f=\rho\omega^{\rho-1}q^{\ast}\Sigma_{\rho}q=\sigma,\quad\sigma=\pm 1.$ (88)
    When $t$ is sufficiently small, there exist a vector $\widetilde{f}(t)=f+O(t^{1/\rho})$ and a real scalar $\widetilde{\omega}(t)=\omega+O(t^{1/\rho})$ satisfying
    $\widetilde{\Theta}_{\rho}(t)\widetilde{f}(t)=\widetilde{\omega}(t)\widetilde{f}(t),\qquad\widetilde{f}^{\ast}(t)\Gamma_{\rho}\widetilde{f}(t)=\sigma,$
    such that the vector
    $\widetilde{u}_{\omega}(t)=\widetilde{U}_{\rho}(t)\widetilde{f}(t)=E_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ G_{\rho}\end{array}\right]q+\sum_{k=1}^{m}U_{k}l_{k\rho}(t)+O(t),$ (89)
    with
    $l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\omega q^{T}&\ldots&(t^{1/\rho}\omega)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
    satisfies
    $(C+tD)\widetilde{u}_{\omega}(t)=(\alpha I+t^{1/\rho}\widetilde{\omega}(t))\widetilde{u}_{\omega}(t),\qquad\widetilde{u}_{\omega}^{\ast}(t)\Delta\widetilde{u}_{\omega}(t)=t^{1-1/\rho}\sigma.$
  2. (b)
    Suppose
    $\Omega=\left[\begin{array}[]{cc}\mu&0\\ 0&\bar{\mu}\end{array}\right],$
    where $\mu,\bar{\mu}$ are nonreal simple eigenvalues of $\Theta_{\rho}(S_{\rho})$. There are vectors
    $f=\left[\begin{array}[]{cccc}q^{T}&\mu q^{T}&\ldots&\mu^{\rho-1}q^{T}\end{array}\right]^{T},\quad f^{(c)}=\left[\begin{array}[]{cccc}(q^{(c)})^{\ast}&\mu(q^{(c)})^{\ast}&\ldots&\mu^{\rho-1}(q^{(c)})^{\ast}\end{array}\right]^{\ast}.$
    that satisfy
    $\Theta_{\rho}(S_{\rho})f=\mu f,\quad\Theta_{\rho}(S_{\rho})f^{(c)}=\bar{\mu}f^{(c)},\quad S_{\rho}q=\mu^{\rho}q,\quad S_{\rho}q^{(c)}=\bar{\mu}^{\rho}q^{(c)},$
    and
    $(f^{(c)})^{\ast}\Gamma_{\rho}f^{(c)}=f^{\ast}\Gamma_{\rho}f=0,\qquad f^{\ast}\Gamma_{\rho}f^{(c)}=\rho(\bar{\mu})^{\rho-1}q^{\ast}\Sigma_{\rho}q^{(c)}=1.$
    For $t$ sufficiently small, there are vectors $\widetilde{f}(t)=f+O(t^{1/\rho})$, $\widetilde{f}^{(c)}(t)=f^{(c)}+O(t^{1/\rho})$, and a nonreal scalar $\widetilde{\mu}(t)=\mu+O(t^{1/\rho})$ satisfying
    $\widetilde{\Theta}_{\rho}(t)\widetilde{f}(t)=\widetilde{\mu}(t)\widetilde{f}(t),\quad\widetilde{\Theta}_{\rho}(t)\widetilde{f}^{(c)}(t)=\overline{\widetilde{\mu}(t)}\widetilde{f}^{(c)}(t),$
    $\widetilde{f}^{\ast}(t)\Gamma_{\rho}\widetilde{f}(t)=(\widetilde{f}^{(c)}(t))^{\ast}\Gamma_{\rho}\widetilde{f}^{(c)}(t)=0,\qquad\widetilde{f}^{\ast}(t)\Gamma_{\rho}\widetilde{f}^{(c)}(t)=1.$
    Then the matrix $\widetilde{U}_{\Omega}(t)=\left[\begin{array}[]{cc}l_{\rho}(t)&l_{\rho}^{(c)}(t)\end{array}\right]$ satisfies
    $(C+tD)\widetilde{U}_{\Omega}(t)=\widetilde{U}_{\Omega}(t)\left[\begin{array}[]{cc}\alpha+t^{1/\rho}\widetilde{\mu}(t)&0\\ 0&\overline{\alpha+t^{1/\rho}\widetilde{\mu}(t)}\end{array}\right],\quad\widetilde{U}_{\Omega}^{\ast}(t)\Delta\widetilde{U}_{\Omega}(t)=t^{1-1/\rho}\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right],$
    where
    $l_{\rho}(t)$ $=$ $\widetilde{U}_{\rho}(t)\widetilde{f}(t)=E_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ G_{\rho}\end{array}\right]q+\sum_{k=1}^{m}U_{k}l_{k\rho}(t)+O(t),$
    $l_{\rho}^{(c)}(t)$ $=$ $\widetilde{U}_{\rho}(t)\widetilde{f}^{(c)}(t)=E_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ G_{\rho}\end{array}\right]q^{(c)}+\sum_{k=1}^{m}U_{k}l_{k\rho}^{(c)}(t)+O(t)$
    with
    $l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\mu q^{T}&\ldots&(t^{1/\rho}\mu)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
    and
    $l_{k\rho}^{(c)}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\mu(q^{(c)})^{\ast}&\ldots&(t^{1/\rho}\mu)^{\rho-1}(q^{(c)})^{\ast}\end{array}\right]^{\ast}&\\ \qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho.\end{array}\right$.
Proof. (a) Since $\omega$ is a simple real eigenvalue of the $\Gamma_{\rho}$-Hermitian matrix $\Theta_{\rho}(S_{\rho})$, following Proposition 3 there is $f\neq 0$ of the form (87) that satisfies (88). Then there is
$\widehat{f}(t)=f+O(t^{1/\rho}),$
such that $\widetilde{\Theta}_{\rho}(t)\widehat{f}(t)=\widetilde{\omega}(t)\widehat{f}(t)$ with $\widetilde{\omega}(t)=\omega+O(t^{1/\rho})$, and for $\widetilde{U}_{\rho}(t)$ given in Theorem 8,
$\widehat{f}^{\ast}(t)\widetilde{U}_{\rho}(t)\Delta\widetilde{U}_{\rho}(t)\widehat{f}(t)=\widehat{f}^{\ast}(t)(t^{1-1/\rho}\Gamma_{\rho})\widehat{f}(t)=t^{1-1/\rho}\widetilde{\sigma},\qquad\widetilde{\sigma}=\sigma+O(t^{1/\rho}).$
Note $\widetilde{\sigma}$ is real. When $t$ is sufficiently small, there is $r=1+O(t^{1/\rho})\in{\mathbb{R}}$ such that $\widetilde{\sigma}r^{2}=\sigma$. By setting $\widetilde{f}(t)=\widehat{f}(t)r=f+O(t^{1/\rho})$, and $\widetilde{u}_{\omega}(t)=\widetilde{U}_{\omega}(t)\widetilde{f}(t)$, one has
$(C+tD)\widetilde{u}_{\omega}(t)=(\alpha+t^{1/\rho}\widetilde{\omega}(t))\widetilde{u}_{\omega}(t),\qquad\widetilde{u}_{\omega}^{\ast}(t)\Delta\widetilde{u}_{\omega}(t)=t^{1-1/\rho}\sigma,$
and $\widetilde{u}_{\omega}(t)$ has the formula as (89). Since $C+tD$ is $\Delta$-Hermitian, the above relations show that $\widetilde{\omega}(t)$ must be real.
(b) Let $f\neq 0$ satisfy
$\Theta_{\rho}(S_{\rho})f=\mu f.$
Because $\Theta_{\rho}(S_{\rho})$ is $\Gamma_{\rho}$-Hermitian and $\mu$ is simple, following Proposition 3, there is $f^{(c)}$ such that both $f,f^{(c)}$ have the given forms, and $F=\left[\begin{array}[]{cc}f&f^{(c)}\end{array}\right]$ satisfies
$\Theta_{\rho}(S_{\rho})F=F\left[\begin{array}[]{cc}\mu&0\\ 0&\bar{\mu}\end{array}\right],\quad F^{\ast}\Gamma_{\rho}F=\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right].$
Similarly, one can show the existence of $\widetilde{f}(t)$, $\widetilde{f}^{(c)}(t)$, and $\widetilde{\mu}(t)$. The results can be derived by setting $\widetilde{U}_{\Omega}(t)=\left[\begin{array}[]{cc}l_{\rho}(t)&l^{(c)}_{\rho}(t)\end{array}\right]=\widetilde{U}_{\rho}(t)\widetilde{F}(t)$, where $\widetilde{F}(t)=\left[\begin{array}[]{cc}\widetilde{f}(t)&\widetilde{f}^{(c)}(t)\end{array}\right]=F+O(t^{1/\rho})$ and satisfies
$\widetilde{\Theta}_{\rho}(t)\widetilde{F}(t)=\widetilde{F}(t)\left[\begin{array}[]{cc}\widetilde{\mu}(t)&0\\ 0&\overline{\widetilde{\mu}(t)}\end{array}\right],\quad\widetilde{\mu}(t)=\mu+O(t^{1/\rho}),\quad\widetilde{F}^{\ast}(t)\Gamma_{\rho}\widetilde{F}(t)=\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right].$
4 Invariant subspace perturbations for Hamiltonian matrices
We now consider the invariant subspace perturbations for Hamiltonian matrices. We assume a Hamiltonian matrix $H$ is perturbed to $H+tT$, where $t$ is sufficiently small and $T$ is another Hamiltonian matrix. We will also focus on the following two cases:
  1. (a)
    a single defective nonimaginary eigenvalue $\lambda$ (paired with $\bar{\lambda}$)
  2. (b)
    a single defective imaginary eigenvalue $\imath\alpha$, where $\alpha\in{\mathbb{R}}$.
Note $H$ is Hamiltonian if and only if $\imath H$ is $\imath J_{n}-$Hermitian. If we set $C=-\imath H$ , $D=-\imath T$, and $\Delta=-\imath J_{n}$, we still have (7), (8), (9). The only difference is that $\Delta$ is congruent to $\mathop{\rm diag}\nolimits(I_{n},-I_{n})$ in this situation. Replacing the nonreal eigenvalue $\lambda$ in (8) by $\imath\lambda$ with a nonimaginary $\lambda$, we have
$C_{T}=-\imath\widehat{H}_{T},\quad\Delta_{T}=U_{T}^{\ast}(-\imath J_{n})U_{T},$
where
$\begin{array}[]{lll}\widehat{H}_{T}=\left[\begin{array}[]{cc}\lambda I_{p}+\imath N&0\\ 0&-(\lambda I_{p}+\imath N)^{\ast}\end{array}\right],&\Delta_{T}=\left[\begin{array}[]{cc}0&I_{p}\\ I_{p}&0\end{array}\right],&\mbox{case (a)}\\ \widehat{H}_{T}=\imath(\alpha I+N),&\Delta_{T}=\mathop{\rm diag}\nolimits(\Gamma_{1},\ldots,\Gamma_{m}),&\mbox{case (b)}\end{array}$
and $\Gamma_{1},\ldots,\Gamma_{m}$ are defined in (9). In terms of $H$ and $J_{n}$, one has
$HU=U\left[\begin{array}[]{cc}\widehat{H}_{T}&0\\ 0&\widehat{H}_{C}\end{array}\right],\quad U^{\ast}J_{n}U=\imath\left[\begin{array}[]{cc}\Delta_{T}&0\\ 0&\Delta_{C}\end{array}\right],\quad U=\left[\begin{array}[]{cc}U_{T}&U_{C}\end{array}\right],$ (92)
where $U$ is the same as (7), $\widehat{H}_{T}$ is defined above, and $\widehat{H}_{C}=\imath C_{C}$.
4.1 Nonimaginary eigenvalue case
Based on the equivalence relations described above, we will use the results developed in Subsection 3.1. In this case, $U_{T}=\left[\begin{array}[]{cc}V&V_{c}\end{array}\right]$ satisfies
$HU_{T}=U_{T}\widehat{H}_{T}=U_{T}\left[\begin{array}[]{cc}\lambda I_{p}+\imath N&0\\ 0&-(\lambda I_{p}+\imath N)^{\ast}\end{array}\right],\qquad U_{T}^{\ast}J_{n}U_{T}=\imath\Delta_{T}=\imath\left[\begin{array}[]{cc}0&I_{p}\\ I_{p}&0\end{array}\right].$
Define the unitary block diagonal matrix
$\widehat{P}=\mathop{\rm diag}\nolimits(\widehat{P}_{1},\ldots,\widehat{P}_{m}),$ (93)
where
$\widehat{P}_{j}=\mathop{\rm diag}\nolimits(I_{s_{j}},(-\imath)I_{s_{j}},\ldots,(-\imath)^{j-1}I_{s_{j}}),\quad j=1,\ldots,m.$
One has
$\widehat{P}^{\ast}(\imath N)\widehat{P}=N.$
Hence
$H_{T}:=\left[\begin{array}[]{cc}\widehat{P}&0\\ 0&-\imath\widehat{P}\end{array}\right]^{\ast}\widehat{H}_{T}\left[\begin{array}[]{cc}\widehat{P}&0\\ 0&-\imath\widehat{P}\end{array}\right]=\left[\begin{array}[]{cc}\lambda I_{p}+N&0\\ 0&-(\lambda I_{p}+N)^{\ast}\end{array}\right],$
and for $\Phi_{T}=\left[\begin{array}[]{cc}\Xi&\Xi_{c}\end{array}\right]$ with
[…]
one has
$H\Phi_{T}=\Phi_{T}H_{T},\qquad\Phi_{T}^{\ast}J_{n}\Phi_{T}=J_{p}.$
Since $\imath J_{n}$ and $\Delta_{T}$ are congruent to $\mathop{\rm diag}\nolimits(I_{n},-I_{n})$ and $\mathop{\rm diag}\nolimits(I_{p},-I_{p})$, respectively, $\Delta_{C}$ in (92) is congruent to $\mathop{\rm diag}\nolimits(I_{n-p},-I_{n-p})$. Therefore, there is a $\widehat{P}_{C}$ such that
$\widehat{P}_{C}^{\ast}(\imath\Delta_{C})\widehat{P}_{C}=J_{n-p}.$ (94)
Let $\Phi_{C}=U_{C}\widehat{P}_{C}$, $\Phi=\left[\begin{array}[]{cc}\Phi_{T}&\Phi_{C}\end{array}\right]$ and $H_{C}=\widehat{P}_{C}^{\ast}\widehat{H}_{C}\widehat{P}_{C}$. Then (92) becomes
$H\Phi=\Phi\left[\begin{array}[]{cc}H_{T}&0\\ 0&H_{C}\end{array}\right],\quad\Phi^{\ast}J_{n}\Phi=\left[\begin{array}[]{cc}J_{p}&0\\ 0&J_{n-p}\end{array}\right].$ (95)
Partition
$\Xi=\left[\begin{array}[]{ccc}\Xi_{1}&\ldots&\Xi_{m}\end{array}\right],\qquad\Xi_{c}=\left[\begin{array}[]{ccc}\Xi_{1}^{(c)}&\ldots&\Xi_{m}^{(c)}\end{array}\right],$ (96)
$\Xi_{j}=\left[\begin{array}[]{ccc}\Xi_{j1}&\ldots&\Xi_{jj}\end{array}\right],\quad\Xi_{j}^{(c)}=\left[\begin{array}[]{ccc}\Xi_{j1}^{(c)}&\ldots&\Xi_{jj}^{(c)}\end{array}\right],\quad j=1,\ldots,m,$
consistent with the block columns of $N$ and $N_{j}$, respectively. Set
$\Upsilon_{k}=\left[\begin{array}[]{ccc}\Xi_{k1}&\ldots&\Xi_{m1}\end{array}\right],\qquad\Upsilon_{k}^{(c)}=\left[\begin{array}[]{ccc}\Xi_{kk}^{(c)}&\ldots&\Xi_{mm}^{(c)}\end{array}\right].$ (97)
Tracing back to the matrix $C+tD=-\imath(H+tT)$, the block $M_{11}$ defined in (44) is
$M_{11}=V_{c}^{\ast}(-J_{n}T)V=(\imath\widehat{P})\Xi_{c}^{\ast}(J_{n}T)\Xi\widehat{P}^{\ast}$
and
$W_{k}=-\left[\begin{array}[]{ccc}(-\imath)^{k}I_{s_{k}}&&\\ &\ddots&\\ &&(-\imath)^{m}I_{s_{m}}\end{array}\right]\Psi_{k},$
where
$\Psi_{k}=(\Upsilon_{k}^{(c)})^{\ast}(J_{n}T)\Upsilon_{k}=\left[\begin{array}[]{ccc}(\Xi_{kk}^{(c)})^{\ast}(J_{n}T)\Xi_{k1}&\ldots&(\Xi_{kk}^{(c)})^{\ast}(J_{n}T)\Xi_{m1}\\ \vdots&\ddots&\vdots\\ (\Xi_{mm}^{(c)})^{\ast}(J_{n}T)\Xi_{k1}&\ldots&(\Xi_{mm}^{(c)})^{\ast}(J_{n}T)\Xi_{m1}\end{array}\right].$ (98)
Then
$G_{\rho}=\widehat{G}_{\rho},\quad G_{\rho}^{(c)}=\left[\begin{array}[]{ccc}-\imath I_{s_{\rho}}&&\\ &\ddots&\\ &&(-\imath)^{m-\rho}I_{s_{m}}\end{array}\right]\widehat{G}^{(c)}_{\rho},\quad S_{\rho}=-(-\imath)^{\rho}\widehat{S}_{\rho}$ (99)
with
$\widehat{G}_{\rho}=-\Psi_{\rho+1}^{-1}(\Upsilon_{\rho+1}^{(c)})^{\ast}J_{n}T\Xi_{\rho 1},\quad\widehat{G}^{(c)}_{\rho}=-\Psi_{\rho+1}^{-\ast}\Upsilon_{\rho+1}^{\ast}J_{n}T\Xi_{\rho\rho}^{(c)},$
$\widehat{S}_{\rho}=\left\{\begin{array}[]{ll}(\Xi_{\rho\rho}^{(c)})^{\ast}J_{n}T(\Xi_{\rho 1}+\Upsilon_{\rho+1}\widehat{G}_{\rho})&\rho<m\\ (\Xi_{mm}^{(c)})^{\ast}J_{n}T\Xi_{m1}&\rho=m.\end{array}\right$. (102)
Based on the results in Subsection 3.1, we have the following results for Hamiltonian matrices.
Theorem 11
Let $H$ and $T$ be Hamiltonian matrices and $H$ satisfy (95) with $\lambda$ being nonimaginary, $\Phi=\left[\begin{array}[]{cc}\Phi_{T}&\Phi_{C}\end{array}\right]$, $\Phi_{T}=\left[\begin{array}[]{cc}\Xi&\Xi_{c}\end{array}\right]$. Suppose $t$ is sufficiently small. Let $\Psi_{1},\ldots,\Psi_{m}$ be defined in (98). For a fixed $\rho\in\{1,\ldots,m\}$ assume $\Psi_{\rho+1}$ is invertible if $\rho<m$, and let $\gamma_{1}^{(\rho)},\ldots,\gamma_{s_{\rho}}^{(\rho)}$ be the $s_{\rho}$ eigenvalues of the matrix $\widehat{S}_{\rho}$, where $\widehat{S}_{\rho}$ is defined in (102).
  1. (a)
[…]
    $\lambda_{ij}^{(\rho)}(t)=\lambda+t^{1/\rho}\mu_{ij}^{(\rho)}+o(t^{1/\rho}),\quad i=1,\ldots,s_{\rho},\quad j=1,\ldots,\rho.$
  2. (b)
    Suppose all $\Psi_{1},\ldots,\Psi_{m}$ are invertible. Let $\Theta_{\rho}(-\widehat{S}_{\rho})$ be defined as (4) with $S_{\rho}$ replaced by $\widehat{S}_{\rho}$. Then for the matrices
    $\widetilde{\Xi}_{\rho}(t)$ $=$ $\Upsilon_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}&0\\ \widehat{G}_{\rho}&0\end{array}\right]+\sum_{k=1}^{m}\Xi_{k}\widetilde{X}_{k\rho}(t)+O(t),$
    $\widetilde{\Xi}_{\rho}^{(c)}(t)$ $=$ $\Upsilon_{\rho}^{(c)}\left[\begin{array}[]{cc}0&I_{s_{\rho}}\\ 0&\widehat{G}_{\rho}^{(c)}\end{array}\right]+\sum_{k=1}^{m}\Xi_{k}^{(c)}\widetilde{X}_{k\rho}^{(c)}(t)+O(t),$
    with
    $\widetilde{X}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \mathop{\rm diag}\nolimits(0_{s_{\rho}},t^{1/\rho}I_{s_{\rho}},\ldots,t^{1-1/\rho}I_{s_{\rho}})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
    and
    $\widetilde{X}^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \mathop{\rm diag}\nolimits(t^{1-1/\rho}I_{s_{\rho}},\ldots,t^{1/\rho}I_{s_{\rho}},0_{s_{\rho}})+O(\left[\begin{array}[]{cccc}t&t^{1-1/\rho}&\ldots&t^{1/\rho}\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho,\end{array}\right$.
    where $\Xi_{k},\Xi_{k}^{(c)}$ and $\Upsilon_{\rho},\Upsilon_{\rho}^{(c)}$ are defined in (96) and (97), respectively, the matrix $\widetilde{\Phi}_{\rho}(t)=\left[\begin{array}[]{cc}\widetilde{\Xi}_{\rho}(t)&\widetilde{\Xi}_{\rho}^{(c)}(t)\end{array}\right]$ satisfies
    $(H+tT)\widetilde{\Phi}_{\rho}(t)=\widetilde{\Phi}_{\rho}(t)\left[\begin{array}[]{cc}\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}^{(h)}(t)&0\\ 0&-(\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}^{(h)}(t))^{\ast}\end{array}\right]$
    and $\widetilde{\Phi}^{\ast}_{\rho}(t)J_{n}\widetilde{\Phi}_{\rho}(t)=t^{1-1/\rho}J_{\rho s_{\rho}}$, where $\widetilde{\Theta}_{\rho}^{(h)}(t)=\Theta_{\rho}(-\widehat{S}_{\rho})+O(t^{1/\rho})$.
Proof. In Theorem 5, set $C=-\imath H$, $D=-\imath T$, $\Delta=-\imath J_{n}$, and replace $\lambda$ by $\imath\lambda$ and $U$ by $\Phi=U\mathop{\rm diag}\nolimits(\widehat{P},-\imath\widehat{P},\widehat{P}_{C})$, where $\widehat{P}$ and $\widehat{P}_{C}$ are defined in (93) and (94). For $t$ sufficiently small, from (51) one has
$(H+tT)\widetilde{\Phi}_{\rho}(t)=\widetilde{\Phi}_{\rho}(t)\left[\begin{array}[]{cc}\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}^{(h)}(t)&0\\ 0&-(\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}^{(h)}(t))^{\ast}\end{array}\right]$
with (by using (99))
$\widetilde{\Theta}_{\rho}^{(h)}(t):=\widehat{P}_{\rho}^{\ast}(\imath\widetilde{\Theta}_{\rho}(S_{\rho}))\widehat{P}_{\rho}=\Theta_{\rho}(\imath^{\rho}S_{\rho})+O(t^{1/\rho})=\Theta_{\rho}(-\widehat{S}_{\rho})+O(t^{1/\rho})$
and
$\widetilde{\Phi}_{\rho}^{\ast}(t)J_{n}\widetilde{\Phi}_{\rho}(t)=t^{1-1/\rho}J_{p},$
where
$\widetilde{\Phi}_{\rho}(t)=\widetilde{U}_{\rho}(t)\left[\begin{array}[]{cc}\widehat{P}_{\rho}&0\\ 0&-\imath\widehat{P}_{\rho}\end{array}\right]=\left[\begin{array}[]{cc}\widetilde{\Xi}_{\rho}(t)&\widetilde{\Xi}_{\rho}^{(c)}(t)\end{array}\right]$
and
$\widetilde{\Xi}_{\rho}(t)$ $=$ $\widetilde{V}_{\rho}(t)\widehat{P}_{\rho}=VX_{\rho}(t)\widehat{P}_{\rho}+O(t)=\Xi\widehat{P}^{\ast}X_{\rho}(t)\widehat{P}_{\rho}+O(t),$
$\widetilde{\Xi}_{\rho}^{(c)}(t)$ $=$ $\widetilde{V}^{(c)}_{\rho}(t)(-\imath\widehat{P}_{\rho})=V_{c}\widehat{X}_{\rho}^{(c)}(t)(-\imath\widehat{P}_{\rho})+O(t)=\Xi_{c}\widehat{P}^{\ast}\widehat{X}_{\rho}^{(c)}(t)\widehat{P}_{\rho}+O(t),$
with $X_{\rho}(t)$ defined in (5) and $\widehat{X}_{\rho}^{(c)}(t)$ in (72). The results can be proven with the above relations and the block forms of $X_{\rho}(t)$ and $\widehat{X}_{\rho}^{(c)}(t)$.
Theorem 12
Under the conditions of Theorem 11, let $F,F^{(c)}\in{\mathbb{C}}^{\rho s_{\rho}\times\ell}$ be full column rank matrices satisfying
$F$ $=$ $\left[\begin{array}[]{cccc}Q^{T}&(Q\Omega)^{T}&\ldots&(Q\Omega^{\rho-1})^{T}\end{array}\right]^{T},\quad-\widehat{S}_{\rho}Q=Q\Omega^{\rho},$
$F^{(c)}$ $=$ $\left[\begin{array}[]{cccc}\Omega^{\rho-1}(Q^{(c)})^{\ast}&\ldots&\Omega(Q^{(c)})^{\ast}&(Q^{(c)})^{\ast}\end{array}\right]^{\ast},\quad-\widehat{S}_{\rho}^{\ast}Q^{(c)}=Q^{(c)}(\Omega^{\ast})^{\rho}$
and
$\Theta_{\rho}(-\widehat{S}_{\rho})F=F\Omega,\quad\Theta_{\rho}^{\ast}(-\widehat{S}_{\rho})F^{(c)}=F^{(c)}\Omega^{\ast},\quad F^{\ast}F^{(c)}=I_{\ell}$
with the condition $\Lambda(\Omega)\cap(\Lambda(\Theta_{\rho}(-\widehat{S}_{\rho}))\backslash\Lambda(\Omega))=\emptyset$. Then for $t$ sufficiently small, there exist matrices $\widetilde{F}(t)=F+O(t^{1/\rho})$, $\widetilde{F}^{(c)}(t)=F^{(c)}+O(t^{1/\rho})$, and $\widetilde{\Omega}(t)=\Omega+O(t^{1/\rho})$ satisfying
$\widetilde{\Theta}_{\rho}^{(h)}(t)\widetilde{F}(t)=\widetilde{F}(t)\widetilde{\Omega}(t),\quad(\widetilde{\Theta}_{\rho}^{(h)}(t))^{\ast}\widetilde{F}^{(c)}(t)=\widetilde{F}^{(c)}(t)\widetilde{\Omega}^{\ast}(t),\qquad\widetilde{F}^{\ast}\widetilde{F}^{(c)}=I_{\ell},$
such that $\widetilde{\Phi}_{\Omega}(t)=\left[\begin{array}[]{cc}L_{\rho}(t)&L_{\rho}^{(c)}(t)\end{array}\right]$ satisfies
$(H+tT)\widetilde{\Phi}_{\Omega}(t)=\widetilde{\Phi}_{\Omega}(t)\left[\begin{array}[]{cc}\lambda I_{\ell}+t^{1/\rho}\widetilde{\Omega}(t)&0\\ 0&-(\lambda I_{\ell}+t^{1/\rho}\widetilde{\Omega}(t))^{\ast}\end{array}\right],\qquad\widetilde{\Phi}^{\ast}_{\Omega}(t)J_{n}\widetilde{\Phi}_{\Omega}(t)=t^{1-1/\rho}J_{\ell},$
where
$L_{\rho}(t)$ $=$ $\widetilde{\Xi}_{\rho}(t)\widetilde{F}(t)=\Upsilon_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ \widehat{G}_{\rho}\end{array}\right]Q+\sum_{k=1}^{m}\Xi_{k}L_{k\rho}(t)+O(t),$
$L_{\rho}^{(c)}(t)$ $=$ $\widetilde{\Xi}_{\rho}^{(c)}(t)\widetilde{F}^{(c)}(t)=\Upsilon_{\rho}^{(c)}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ \widehat{G}_{\rho}^{(c)}\end{array}\right]Q^{(c)}+\sum_{k=1}^{m}\Xi_{k}^{(c)}L_{k\rho}^{(c)}(t)+O(t)$
with
$L_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}&\\ \qquad\qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
and
$L^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\Omega)^{\rho-1}(Q^{(c)})^{\ast}&\ldots&(t^{1/\rho}\Omega)(Q^{(c)})^{\ast}&0\end{array}\right]^{\ast}&\\ \qquad\qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho.\end{array}\right$.
Proof. The proof is essentially the same as that for Theorem 6.
Corollary 13
Suppose $\omega$ is a simple eigenvalue of $\Theta_{\rho}(-\widehat{S}_{\rho})$, and the vectors
$f=\left[\begin{array}[]{cccc}q^{T}&\omega q^{T}&\ldots&\omega^{\rho-1}q^{T}\end{array}\right]^{T},\quad f^{(c)}=\left[\begin{array}[]{cccc}\omega^{\rho-1}(q^{(c)})^{\ast}&\ldots&\omega(q^{(c)})^{\ast}&(q^{(c)})^{\ast}\end{array}\right]^{\ast}$
satisfy
$\Theta_{\rho}(-\widehat{S}_{\rho})f=\omega f,\quad\Theta_{\rho}^{\ast}(-\widehat{S}_{\rho})f^{(c)}=\bar{\omega}f^{(c)},\quad f^{\ast}f^{(c)}=\rho\bar{\omega}^{\rho-1}q^{\ast}q^{(c)}=1.$
For $t$ sufficiently small, there are vectors
$\widetilde{f}(t)=f+O(t^{1/\rho}),\quad\widetilde{f}^{(c)}(t)=f^{(c)}+O(t^{1/\rho}),$
and a scalar $\widetilde{\omega}(t)=\omega+O(t^{1/\rho})$ satisfying
$\widetilde{\Theta}_{\rho}^{(h)}(t)\widetilde{f}(t)=\widetilde{\omega}(t)\widetilde{f}(t),\qquad(\widetilde{\Theta}_{\rho}^{(h)}(t))^{\ast}\widetilde{f}^{(c)}(t)=\overline{\widetilde{\omega}(t)}\widetilde{f}^{(c)}(t),\quad\widetilde{f}^{\ast}(t)\widetilde{f}^{(c)}(t)=1,$
such that $\widetilde{\Phi}_{\omega}(t)=\left[\begin{array}[]{cc}l_{\rho}(t)&l_{\rho}^{(c)}(t)\end{array}\right]$ satisfies
$(H+tT)\widetilde{\Phi}_{\omega}(t)=\widetilde{\Phi}_{\omega}(t)\left[\begin{array}[]{cc}\lambda+t^{1/\rho}\widetilde{\omega}(t)&0\\ 0&-\overline{(\lambda+t^{1/\rho}\widetilde{\omega}(t))}\end{array}\right],\qquad\widetilde{\Phi}_{\omega}^{\ast}(t)J_{n}\widetilde{\Phi}_{\omega}(t)=t^{1-1/\rho}J_{1},$
and
$l_{\rho}(t)$ $=$ $\widetilde{\Xi}_{\rho}(t)\widetilde{f}(t)=\Upsilon_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ \widehat{G}_{\rho}\end{array}\right]q+\sum_{k=1}^{m}\Xi_{k}l_{k\rho}(t)+O(t),$
$l_{\rho}^{(c)}(t)$ $=$ $\widetilde{\Xi}_{\rho}^{(c)}(t)\widetilde{f}^{(c)}(t)=\Upsilon_{\rho}^{(c)}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ \widehat{G}_{\rho}^{(c)}\end{array}\right]q^{(c)}+\sum_{k=1}^{m}\Xi_{k}^{(c)}l_{k\rho}^{(c)}(t)+O(t)$
with
$l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\omega q^{T}&\ldots&(t^{1/\rho}\omega)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho,\end{array}\right$.
$l^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\omega)^{\rho-1}(q^{(c)})^{\ast}&\ldots&t^{1/\rho}\omega(q^{(c)})^{\ast}&0\end{array}\right]^{\ast}+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho.\end{array}\right$.
4.2 Purely imaginary eigenvalue case
The results will be derived directly from those in Subsection 3.2 with $C=-\imath H$, $D=-\imath T$ and $\Delta=-\imath J_{n}$. For consistency, we replace $U=\left[\begin{array}[]{cc}U_{T}&U_{C}\end{array}\right]$ with $\Phi=\left[\begin{array}[]{cc}\Phi_{T}&\Phi_{C}\end{array}\right]$. Partition
$\Phi_{T}=\left[\begin{array}[]{cccc}\Phi_{1}&\Phi_{2}&\ldots&\Phi_{m}\end{array}\right],$ (111)
$\Phi_{j}=\left[\begin{array}[]{cccc}\Phi_{j1}&\Phi_{j2}&\ldots&\Phi_{jj}\end{array}\right],\quad j=1,\ldots,m,$
that are consistent with the block columns of $N$ and $N_{j}$, respectively, and set
$\Upsilon_{\rho}=\left[\begin{array}[]{ccc}\Phi_{\rho 1}&\ldots&\Phi_{m1}\end{array}\right]$ (112)
for $\rho\in\{1,\ldots,m\}$. We still have (92), and in this case $\widehat{H}_{T}=\imath(\alpha I_{p}+N)$ with $\alpha\in{\mathbb{R}}$, and $\Delta_{T}$ is defined in (9). With the new notations, the matrix $W_{k}$ corresponding to $C$ defined in (82) becomes
$W_{k}=-\left[\begin{array}[]{ccc}\Sigma_{k}&&\\ &\ddots&\\ &&\Sigma_{m}\end{array}\right]\Psi_{k},\qquad\Psi_{k}=\Upsilon_{k}^{\ast}(J_{n}T)\Upsilon_{k}=\Psi_{k}^{\ast}.$ (113)
Hence
$G_{\rho}$ $=$ $-\Psi_{\rho+1}^{-1}\Upsilon_{\rho+1}^{\ast}(J_{n}T)\Phi_{\rho 1},$
$S_{\rho}$ $=$ $-\Sigma_{\rho}\widehat{S}_{\rho},\qquad\widehat{S}_{\rho}=\widehat{S}_{\rho}^{\ast}=\left\{\begin{array}[]{ll}\Phi_{\rho 1}^{\ast}(J_{n}T)(\Phi_{\rho 1}+\Upsilon_{\rho+1}G_{\rho})&\rho<m\\ \Phi_{m1}^{\ast}(J_{n}T)\Phi_{m1}&\rho=m.\end{array}\right$. (117)
Theorem 14
Suppose both $H$ and $T$ are Hamiltonian matrices. For a given $\rho\in\{1,\ldots,m\}$, assume $\Psi_{\rho+1}$ defined in (113) is invertible if $\rho<m$. Let $G_{\rho}$ and $S_{\rho}$ be defined in (LABEL:defhgs) and $\gamma_{1}^{(\rho)},\ldots,\gamma_{s_{\rho}}^{(\rho)}$ be the $s_{\rho}$ eigenvalues of the matrix $S_{\rho}$.
  1. (a)
[…]
    $\lambda_{ij}^{(\rho)}(t)=\imath(\alpha+t^{1/\rho}\mu_{ij}^{(\rho)})+o(t^{1/\rho}),\qquad i=1,\ldots,s_{\rho},\quad j=1,\ldots,\rho.$
  2. (b)
    Suppose all $\Psi_{1},\ldots,\Psi_{m}$ are invertible. Let $\Theta_{\rho}(S_{\rho})$ be defined as (4) with $S_{\rho}$ given in (LABEL:defhgs). Then for $t$ sufficiently small, there is a matrix
    $\widetilde{\Phi}_{\rho}(t)=\Upsilon_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}&0\\ G_{\rho}&0\end{array}\right]+\sum_{k=1}^{m}\Phi_{k}\widetilde{X}_{k\rho}(t)+O(t),\\$
    with $G_{\rho}$ given in (LABEL:defhgs) and
    $\widetilde{X}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \mathop{\rm diag}\nolimits(0_{s_{\rho}},t^{1/\rho}I_{s_{\rho}},\ldots,t^{1-1/\rho}I_{s_{\rho}})+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
    where $\Phi_{1},\ldots,\Phi_{m}$ are the block columns of $\Phi_{T}$ defined in (111) and $\Upsilon_{\rho}$ is defined in (112), such that
    $(H+tT)\widetilde{\Phi}_{\rho}(t)=\widetilde{\Phi}_{\rho}(t)(\imath(\alpha I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}^{(h)}(t))),\qquad\widetilde{\Theta}_{\rho}^{(h)}(t)=\Theta_{\rho}(S_{\rho})+O(t^{1/\rho})$
    and $\widetilde{\Phi}_{\rho}^{\ast}(t)J_{n}\widetilde{\Phi}_{\rho}(t)=t^{1-1/\rho}(\imath\Gamma_{\rho})$.
Proof. It is directly from Theorem 8.
Theorem 15
Under the conditions of Theorem 14, suppose $F\in{\mathbb{C}}^{\rho s_{\rho}\times\ell}$ is a full column rank matrix of the form (84) that satisfies (83), where $\Lambda(\Omega)\subseteq\Lambda(\Theta_{\rho}(S_{\rho}))$. Assume the eigenvalues of $\Omega$ are either real or in complex conjugate pairs, and $\Lambda(\Omega)\cap(\Lambda(\Theta_{\rho}(S_{\rho}))\backslash\Lambda(\Omega))=\emptyset$. Then for $t$ sufficiently small, there is a matrix of the form
$\widetilde{\Phi}_{\Omega}(t)=\Upsilon_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ G_{\rho}\end{array}\right]Q+\sum_{k=1}^{m}\Phi_{k}L_{k\rho}(t)+O(t)$
with
$L_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
such that
$(H+tT)\widetilde{\Phi}_{\Omega}(t)=\widetilde{\Phi}_{\Omega}(t)(\imath(\alpha I+t^{1/\rho}\widetilde{\Omega}(t))),\qquad\widetilde{\Omega}(t)=\Omega+O(t^{1/\rho}),$
and $\widetilde{\Phi}^{\ast}_{\Omega}(t)J_{n}\widetilde{\Phi}_{\Omega}(t)=t^{1-1/\rho}(\imath\Delta_{\Omega})$ with $\Delta_{\Omega}$ given in (83).
Proof. It is directly from Theorem 9.
Corollary 16
  1. (a)
    Suppose $\omega$ is a simple real eigenvalue of $\Theta_{\rho}(S_{\rho})$. There is a vector
    $f=\left[\begin{array}[]{cccc}q^{T}&\omega q^{T}&\ldots&\omega^{\rho-1}q^{T}\end{array}\right]^{T}$
    that satisfies
    $\Theta_{\rho}(S_{\rho})f=\omega f,\quad S_{\rho}q=\omega^{\rho}q,\quad f^{\ast}\Gamma_{\rho}f=\rho\omega^{\rho-1}q^{\ast}\Sigma_{\rho}q=\sigma,\quad\sigma=\pm 1.$
    When $t$ is sufficiently small, there is a vector $\widetilde{f}(t)=f+O(t^{1/\rho})$ and a real scalar $\widetilde{\omega}(t)=\omega+O(t^{1/\rho})$ satisfying
    $\widetilde{\Theta}_{\rho}^{(h)}(t)\widetilde{f}(t)=\widetilde{\omega}(t)\widetilde{f}(t),\qquad\widetilde{f}^{\ast}(t)\Gamma_{\rho}\widetilde{f}(t)=\sigma,$
    and the vector
    $\widetilde{\phi}_{\omega}(t)=\widetilde{\Phi}_{\rho}(t)\widetilde{f}(t)=\Upsilon_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ G_{\rho}\end{array}\right]q+\sum_{k=1}^{m}\Phi_{k}l_{k\rho}(t)+O(t),$
    with
    $l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\omega q^{T}&\ldots&(t^{1/\rho}\omega)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
    satisfies
    $(H+tT)\widetilde{\phi}_{\omega}(t)=\imath(\alpha I+t^{1/\rho}\widetilde{\omega}(t))\widetilde{\phi}_{\omega}(t),\quad\widetilde{\phi}_{\omega}^{\ast}(t)J_{n}\widetilde{\phi}_{\omega}(t)=t^{1-1/\rho}\imath\sigma.$
  2. (b)
    Suppose
    $\Omega=\left[\begin{array}[]{cc}\mu&0\\ 0&\bar{\mu}\end{array}\right],$
    where $\mu,\bar{\mu}$ are nonreal simple eigenvalues of $\Theta_{\rho}(S_{\rho})$. There are vectors
    $f=\left[\begin{array}[]{cccc}q^{T}&\mu q^{T}&\ldots&\mu^{\rho-1}q^{T}\end{array}\right]^{T},\quad f^{(c)}=\left[\begin{array}[]{cccc}(q^{(c)})^{\ast}&\mu(q^{(c)})^{\ast}&\ldots&\mu^{\rho-1}(q^{(c)})^{\ast}\end{array}\right]^{\ast}$
    that satisfy
    $\Theta_{\rho}(S_{\rho})f=\mu f,\quad S_{\rho}q=\mu^{\rho}q,\quad\Theta_{\rho}(S_{\rho})f^{(c)}=\bar{\mu}f^{(c)},\quad S_{\rho}q^{(c)}=\bar{\mu}^{\rho}q^{(c)},$
    and
    $(f^{(c)})^{\ast}\Gamma_{\rho}f^{(c)}=f^{\ast}\Gamma_{\rho}f=0,\qquad f^{\ast}\Gamma_{\rho}f^{(c)}=\rho\bar{u}^{\rho-1}q^{\ast}\Sigma_{\rho}q^{(c)}=1.$
    When $t$ is sufficiently small, there are vectors $\widetilde{f}(t)=f+O(t^{(1/\rho})$, $\widetilde{f}^{(c)}(t)=f^{(c)}+O(t^{(1/\rho})$, and a nonreal scalar $\widetilde{\mu}(t)=\mu+O(t^{(1/\rho})$ satisfying
    $\widetilde{\Theta}_{\rho}^{(h)}(t)\widetilde{f}(t)=\widetilde{\mu}(t)\widetilde{f}(t),\qquad\widetilde{\Theta}_{\rho}^{(h)}(t)\widetilde{f}^{(c)}(t)=\overline{\widetilde{\mu}(t)}\widetilde{f}^{(c)}(t),$
    $\widetilde{f}^{\ast}(t)\Gamma_{\rho}\widetilde{f}(t)=(\widetilde{f}^{(c)}(t))^{\ast}\Gamma_{\rho}\widetilde{f}^{(c)}(t)=0,\quad\widetilde{f}^{\ast}(t)\Gamma_{\rho}\widetilde{f}^{(c)}(t)=1,$
    such that $\widetilde{\Phi}_{\Omega}(t)=\left[\begin{array}[]{cc}l_{\rho}(t)&l_{\rho}^{(c)}(t)\end{array}\right]$ satisfies
    $(H+tT)\widetilde{\Phi}_{\Omega}(t)$ $=$ $\widetilde{\Phi}_{\Omega}(t)\left[\begin{array}[]{cc}\imath(\alpha+t^{1/\rho}\widetilde{\mu}(t))&0\\ 0&-\,\overline{[\imath(\alpha+t^{1/\rho}\widetilde{\mu}(t))]}\end{array}\right],$
    $\widetilde{\Phi}_{\Omega}^{\ast}(t)J_{n}\widetilde{\Phi}_{\Omega}(t)$ $=$ $\imath t^{1-1/\rho}\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right],$
    where
    $l_{\rho}(t)$ $=$ $\widetilde{\Phi}_{\rho}(t)\widetilde{f}(t)=\Upsilon_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ G_{\rho}\end{array}\right]q+\sum_{k=1}^{m}\Phi_{k}l_{k\rho}(t)+O(t),$
    $l_{\rho}^{(c)}(t)$ $=$ $\widetilde{\Phi}_{\rho}(t)\widetilde{f}^{(c)}(t)=\Upsilon_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ G_{\rho}\end{array}\right]q^{(c)}+\sum_{k=1}^{m}\Phi_{k}l_{k\rho}^{(c)}(t)+O(t)$
    with
    $l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\mu q^{T}&\ldots&(t^{1/\rho}\mu)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
    and
    $l_{k\rho}^{(c)}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\mu)^{\rho-1}(q^{(c)})^{\ast}&\ldots&t^{1/\rho}\omega(q^{(c)})^{\ast}&0\end{array}\right]^{\ast}&\\ \qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho.\end{array}\right$.
4.3 A special case
For Hamiltonian matrices an important question is under what conditions the eigenvalues will stay on the imaginary axis under (small) Hamiltonian perturbations and how to remove them from the imaginary axis with Hamiltonian perturbations. Based on the results established above (and Remark 4), when $t$ is sufficiently small, it is obvious that necessary conditions for $H+tT$ to have imaginary eigenvalues include
  1. 1.
    $H$ has imaginary eigenvalues
  2. 2.
    For an imaginary eigenvalue of $H$, the corresponding $S_{\rho}=-\Sigma_{\rho}\widehat{S}_{\rho}$ defined in (LABEL:defhgs) should have real roots when $\rho$ is odd, or positive real roots when $\rho$ is even.
[…]
In [22] a special case was considered with
$H=\widehat{U}\left[\begin{array}[]{cc|cc}\imath\alpha I_{p}+N&0&G&0\\ 0&F_{2}&0&G_{2}\\ \hline\cr 0&0&-(\imath\alpha I_{p}+N)^{\ast}&0\\ 0&R_{2}&0&-F_{2}^{\ast}\end{array}\right]\widehat{U}^{-1},\quad T=J_{n}K,$
where $\alpha\in{\mathbb{R}}$, $\widehat{U}$ is symplectic ($\widehat{U}^{\ast}J_{n}\widehat{U}=J_{n}$), $H_{C}:=\left[\begin{array}[]{cc}F_{2}&G_{2}\\ R_{2}&-F_{2}^{\ast}\end{array}\right]$ is Hamiltonian and $\imath\alpha\notin\Lambda(H_{C})$. $K$ is positive semidefinite, NN is defined in (1), and
$G=\mathop{\rm diag}\nolimits(G_{1},\ldots,G_{m}),\quad G_{k}=\mathop{\rm diag}\nolimits(0,\ldots,0,I_{s_{k}}),\quad k=1,\ldots,m.$
After a simple two-sided block permutation, we have a nonsingular matrix $U=\left[\begin{array}[]{cc}U_{T}&U_{C}\end{array}\right]$ such that
$HU=U\left[\begin{array}[]{cc}H_{T}&0\\ 0&H_{C}\end{array}\right],\quad U^{\ast}J_{n}U=\left[\begin{array}[]{cc}\widehat{J}_{p}&0\\ 0&J_{n-p}\end{array}\right],$
where
$H_{T}=\mathop{\rm diag}\nolimits(\imath\alpha I_{2s_{1}}+\widehat{H}_{2},\imath\alpha I_{4s_{2}}+\widehat{H}_{4},\ldots,\imath\alpha I_{2ms_{m}}+\widehat{H}_{2m}),\quad\widehat{J}_{p}=\mathop{\rm diag}\nolimits(J_{s_{1}},J_{2s_{2}},\ldots,J_{ms_{m}})$
with
$\widehat{H}_{2k}=\left[\begin{array}[]{cc}N_{k}&G_{k}\\ 0&-N_{k}^{\ast}\end{array}\right],\quad k=1,\ldots,m.$
Let $\widehat{P}_{2k}$ be defined in (93) and
$\Pi_{T}=\mathop{\rm diag}\nolimits(\Pi_{1},\ldots,\Pi_{m}),$
where
[…]
Then
$\Pi_{T}^{\ast}H_{T}\Pi_{T}=\imath(\alpha I+\widehat{N}),\quad\Pi_{T}^{\ast}\widehat{J}_{p}\Pi_{T}=\imath\widehat{\Delta}_{T},$
where
$\widehat{N}=\mathop{\rm diag}\nolimits(\widehat{N}_{2},\widehat{N}_{4},\ldots,\widehat{N}_{2m}),\quad\widehat{\Delta}_{T}=\imath\mathop{\rm diag}\nolimits(\Gamma_{2},\Gamma_{4},\ldots,\Gamma_{2m})$
with
[…]
Note in this special case the sizes of the Jordan blocks of $\imath\alpha$ are all even. Let
$\Phi=\left[\begin{array}[]{cc}U_{T}&U_{C}\end{array}\right]\mathop{\rm diag}\nolimits(\Pi_{T},I)=\left[\begin{array}[]{cc}\Phi_{T}&\Phi_{C}\end{array}\right],$
and partition
$\Phi_{T}=\left[\begin{array}[]{cccc}\Phi_{2}&\Phi_{4}&\ldots&\Phi_{2m}\end{array}\right],\quad\Phi_{2j}=\left[\begin{array}[]{cccc}\Phi_{2j,1}&\Phi_{2j,2}&\ldots&\Phi_{2j,2j}\end{array}\right]$
conformably with the block columns of $\widehat{N}$ and $\widehat{N}_{2k}$. Let
$\Upsilon_{2k}=\left[\begin{array}[]{cccc}\Phi_{2k,1}&\Phi_{2(k+1),1}&\ldots&\Phi_{2m,1}\end{array}\right].$
Since in this case, $\Sigma_{2k}=I_{s_{k}}$ for $k=1,\ldots,m$ and $T=J_{n}K$, following (113) one has
$W_{2k}=-\Upsilon_{2k}^{\ast}(J_{n}(J_{n}K))\Upsilon_{2k}^{\ast}=\Upsilon_{2k}^{\ast}K\Upsilon_{2k}.$
Since $K\geq 0$, the nonsingularity requirement becomes $W_{2k}>0$ for $k=1,\ldots.m$. Then $S_{2\rho}>0$ for any $\rho\in\{1,\ldots.m\}$. Let $\gamma_{j}>0$ be a semisimple eigenvalue of $S_{2\rho}$ with algebraic (geometric) multiplicity $\ell$. Then $\Theta_{2\rho}(S_{2\rho})$ has exactly two real semisimple eigenvalues corresponding to $\gamma_{j}$ with the same algebraic multiplicity $\ell$, denoted by $\mu_{j,\pm}=\pm\gamma_{j}^{1/(2\rho)}$. Consider $\mu_{j,+}$. Let $Q$ be a basis matrix of the eigenvector space of $S_{2\rho}$ corresponding to $\gamma_{j}$, i.e.,
$S_{2\rho}Q=\gamma_{j}Q.$
Then
$F=\left[\begin{array}[]{cccc}Q^{T}&\mu_{j,+}Q^{T}&\ldots&\mu_{j,+}^{2\rho-1}Q^{T}\end{array}\right]^{T}$
has full column rank and satisfies
$\Theta_{2\rho}(S_{2\rho})F=F\Omega,\qquad\Omega=\mu_{j,+}I_{\ell}.$
This shows that $\mu_{j,+}$ is semisimple and $F$ is a basis matrix of the corresponding eigenvector space of $\Theta_{2\rho}(S_{2\rho})$. Because
$\Delta_{\Omega}:=F^{\ast}\Gamma_{2\rho}F=2\rho\mu_{j,+}^{2\rho-1}Q^{\ast}Q>0.$
One can always choose $Q$ such that $\Delta_{\Omega}=I_{\ell}$. When $t>0$ is sufficiently small, following Theorem 15 there are $\widetilde{\Phi}_{\Omega}(t)$ and
$\widetilde{\Omega}(t)=\Omega+O(t^{1/(2\rho)})=\mu_{j,+}I_{\ell}+O(t^{1/(2\rho)})$
such that
$(H+tJ_{n}K)\widetilde{\Phi}_{\Omega}(t)=\widetilde{\Phi}_{\Omega}(t)(\imath(\alpha I_{\ell}+t^{1/(2\rho)}\widetilde{\Omega}(t))),\quad\widetilde{\Phi}_{\Omega}^{\ast}(t)J_{n}\widetilde{\Phi}_{\Omega}(t)=\imath t^{1-1/(2\rho)}I_{\ell}.$
Notice
$\widetilde{\Phi}_{\Omega}^{\ast}(t)J_{n}(H+tJ_{n}K)\widetilde{\Phi}_{\Omega}(t)=-t^{1-1/(2\rho)}(\alpha I_{\ell}+t^{1/(2\rho)}\widetilde{\Omega}(t)).$
Since the left-hand side is Hermitian, so is $\widetilde{\Omega}(t)$. This implies that $H+tJ_{n}K$ has $\ell$ (semi)simple purely imaginary eigenvalues of the form $\imath(\alpha+t^{1/(2\rho)}\mu_{j,+}+o(t^{1/(2\rho)}))$ and the corresponding structure inertia indices are all $1$. In the same way, one can show that $H+tJ_{n}K$ has $\ell$ purely imaginary eigenvalues of the form $\imath(\alpha+t^{1/(2\rho)}\mu_{j,-}+o(t^{1/(2\rho)}))$ and the corresponding structure inertia indices are all 1-1. The same result was derived in [22], which now can be considered as a special case of Theorem 15.
5 Conclusions
By using the recently developed perturbation theory in [28], we have derived structured invariant subspace perturbation results for both the $\Delta$-Hermitian matrices and Hamiltonian matrices corresponding to defective eigenvalues. The results exhibit the fractional orders and symmetric structures of the perturbed eigenvalues and invariant subspaces, and provide the information that are not contained in the standard perturbation theory. The perturbation results only apply to the case where all the blocks $W_{1},\ldots,W_{m}$ are invertible. Currently, it is not clear how to extend the results to general case.
Acknowledgement The author thanks the anonymous referee and the handling editor for their comments and suggestions.